### v1

# Controlling discrete semilinear wave equations toward flocks

Abstract.
In this work, we initiate the research on controlling nonlinear waves propagating on lattices from a completely new perspective. We consider nonlinear waves on a lattice as a system of interacting particles and study their collective flocking behavior. By designing suitable feedback controls, we show that any admissible flock can be reached within a finite amount of time. Finally, we highlight the connection between our flocking problem and a minimal-time problem in the framework of nonlinear Hamilton–Jacobi equations and optimal control theory.
Key words and phrases:
discrete semilinear wave equations; flocks; controls; optimal control; Hamilton–Jacobi equations
2010 Mathematics Subject Classification:
35B65, 35F21, 35Q99, 44A10, 45J05
The work of SS is partially supported by … The work of HVT is partially supported by NSF grant DMS-2348305. The work of MBT is partially supported by NSF Grants CAREER DMS-2303146, DMS-2204795, DMS-2306379.
1. Introduction
1.1. Brief overview
In recent years, many efforts have been made to understand atomic vibrations in anharmonic chains and lattices. Atomic vibrations in anharmonic chains (one-dimensional systems) and lattices (multi-dimensional systems) can be described via discrete wave-type equations [5, 9, 11].
Control theory, which has been extensively studied since the last 60 years, has provided insight into several fields, including finance, medicine, mechanics, engineering, and climatology. Over several years, control theory has significantly shaped the study of wave equations [14].
The goal of our paper is to initiate the research on controlling atomic vibrations in anharmonic chains and lattices via the associated nonlinear discrete wave equations. In particular, nonlinear discrete wave equations can be applied to lattices in solid mechanics such as Peierls’ lattice [9] or to vibrating strings of oscillators like the Fermi-Pasta-UIam-Tsingou lattice [5]. We note that these lattices have a fixed grid size. It is well-known that wave vibrations on lattices can exhibit complicated behaviors. For example, the experiment performed by Fermi, Pasta, Ulam, and Tsingou reported in [5] produced nonintuitive results and now motivates the Fermi-Pasta-Ulam-Tsingou paradox. We aim to design controls to stabilize these physically relevant fixed grid size systems by driving them towards flocks (see Definition 1).
Although numerous collective behavior models have been introduced and investigated to analyze the emergent flocking of self-organized systems, starting with the pioneering work of Cucker–Smale [2, 3, 8], to the best of our knowledge, this is the first time in the literature that the problem of control around flocks for discrete partial differential equations has been addressed and studied. In recent years, there have been important advances in the control theory for kinetic models, particularly the control to flocking of the Cucker–Smale model [1, 10], which serves as our source of inspiration for studying the problem of controlling the propagation of lattice/chains nonlinear waves.
Controlling discrete nonlinear wave equations for numerical purposes has been well studied in the literature [4, 15]. However, the goal of this research is to design a control that is a good approximation for the continuous problem. Thus, the goal is to observe the behavior of the control as the mesh size tends to zero. Since our problem is meant to be applied to lattices that have a fixed size, there are large differences in the theory. The stochastic control problem for discrete linear wave equations with a fixed mesh size was recently studied in [6, 7]. However, to our knowledge, this work will be the first in the control theory of nonlinear partial differential equations to discuss the fixed mesh size approach for a discrete nonlinear wave equation.
1.2. Settings and main goals
Let $D,n\in\mathbb{N}$ be given natural numbers, and let $h=1/D$ be the fixed grid size. Denote by
$\Lambda=\left\{\frac{j}{D}\,:\,j\in\mathbb{Z}\right\}^{n}=\left\{jh\,:\,j\in\mathbb{Z}\right\}^{n}$
the discrete grid in $\mathbb{R}^{n}$.
We consider the following discrete semilinear Klein–Gordon equation with a cubic nonlinearity and a control term
$\begin{cases}w_{tt}-\Delta_{D}w=(\alpha-\beta|w_{t}|^{2})w_{t}+u(l,t)\qquad&\text{ in }\Lambda\times(0,\infty),\\ w(\cdot,0)=w_{0},~{}w_{t}(\cdot,0)=w_{1}\qquad&\text{ on }\Lambda.\end{cases}$ (1.1)
Note that both the semilinear Klein–Gordon equation with cubic and quadratic nonlinearities can be used to describe atomic vibrations in anharmonic lattices [12]. In this work, we only focus on the Klein–Gordon equation with a cubic nonlinearity, while the other case can be treated similarly.
Here, $\alpha,\beta>0$ are given constants. And $w_{0},w_{1}$ are the given initial data, which are $\mathbb{Z}^{n}$-periodic. For $l\in\Lambda$, we write $l^{\prime}\sim l$ if $l^{\prime}-l=\pm he_{k}$ for some $k\in\{1,\ldots,n\}$. The corresponding discrete Laplacian is
$\Delta_{D}w(l,t)=\Delta w(l,t)=\sum_{l^{\prime}\sim l}\frac{w(l^{\prime},t)-w(l,t)}{h^{2}}.$
When there is no confusion, we write $\Delta_{D}w(l,t)=\Delta w(l,t)$ for simplicity. The function $u:\Lambda\times[0,\infty)\to\mathbb{R}$, which is $\mathbb{Z}^{n}$-periodic in $l$, is our control, and we typically assume that
$u\in\mathcal{A}_{M}=\{f\in L^{\infty}(\Lambda\times[0,\infty))\,:\,f\text{ is $\mathbb{Z}^{n}$-periodic in $l$, }\|f\|_{L^{\infty}}\leq M\}$
where $M>0$ is a given positive constant (to be specified later).
As we are in the spatial $\mathbb{Z}^{n}$-periodic setting, we are concerned with the spatial $\mathbb{Z}^{n}$-periodic solution $w$ to (1.1). In particular, our problem can be phrased in the discrete torus
$\overline{\Lambda}=\left\{0,\frac{1}{D},\ldots,\frac{D-1}{D}\right\}^{n},$
where we identify $k+j/D\equiv j/D$ for $k,j\in\mathbb{Z}$. If a function $f:\Lambda\to\mathbb{R}$ is $\mathbb{Z}^{n}$-periodic, we can think of $f$ as a function from $\overline{\Lambda}$ to $\mathbb{R}$ as well, and vice versa. As explained in the introduction, our mesh size $h$ is fixed, which is the same as the lattice considered by Peierls [9].
In this paper, we are interested in using the control $u$ to drive the solution $w$ of (1.1) to a flock where the group velocity is constant. Looking at (1.1), it is clear that the nonlinear term vanishes if
$w_{t}=\bar{v}=\pm\sqrt{\frac{\alpha}{\beta}}\qquad\text{ or }\qquad w_{t}=\bar{v}=0.$
Further, by using the ansatz $w(l,t)=\bar{w}(l)+\bar{v}t$ in (1.1), we see that
$|-\Delta\bar{w}(l)|=|u(l,t)|\leq M\quad\text{ for }l\in\overline{\Lambda},$ (1.2)
which gives a compatibility condition.
Definition 1.
We say that $\varphi:\overline{\Lambda}\times\mathbb{R}\to\mathbb{R}$ is a moving flock if
$\varphi(l,t)=\bar{\varphi}(l)+\bar{v}t\quad\text{ for }(l,t)\in\overline{\Lambda}\times\mathbb{R},$
where $\bar{v}=\pm\sqrt{\frac{\alpha}{\beta}}$ and $\bar{\varphi}$ satisfies (1.2).
We say that $\varphi:\overline{\Lambda}\times\mathbb{R}\to\mathbb{R}$ is a stationary flock if
$\varphi(l,t)=\bar{\varphi}(l)\quad\text{ for }(l,t)\in\overline{\Lambda}\times\mathbb{R},$
where $\bar{\varphi}$ satisfies (1.2).
By symmetry, we only need to consider the moving flocks with $\bar{v}=\sqrt{\alpha/\beta}$. The main questions of interest of our paper are as follows.
  • (1)
    Let $\bar{v}=\sqrt{\alpha/\beta}$ and $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Can we find a control $u\in\mathcal{A}_{M}$ such that $w=\varphi_{1}$ for all $t\geq T$, where $T>0$ is sufficiently large and $\varphi_{1}$ is a moving flock?
  • (2)
    Let $\bar{v}=\sqrt{\alpha/\beta}$ and $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Take an arbitrary moving flock $\varphi_{2}(l,t)=\bar{\varphi}_{2}(l)+\bar{v}t$. Can we find a control $u\in\mathcal{A}_{M}$ such that $w-\varphi_{2}$ is constant for all $t\geq T$ where $T>0$ is sufficiently large?
  • (3)
    Let $\bar{v}=0$ and $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Take an arbitrary stationary flock $\varphi_{3}(l,t)=\bar{\varphi}_{3}(l)$. Can we find a control $u\in\mathcal{A}_{M}$ such that $w-\varphi_{3}$ is constant for all $t\geq T$ where $T>0$ is sufficiently large?
1.3. Main results
Write $x_{l}(t)=w(l,t)$ and $v_{l}(t)=w_{t}(l,t)$ for $l\in\Lambda$ and $t\geq 0$. Then, (1.1) becomes a system of ODEs, for $(l,t)\in\Lambda\times(0,\infty)$,
$\begin{cases}\dot{x}_{l}(t)=v_{l}(t),\\ \displaystyle\dot{v}_{l}(t)=\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}+(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)+u_{l}(t),\end{cases}$ (1.3)
with given initial condition
$\begin{cases}x_{l}(0)=w_{0}(l),\\ v_{l}(0)=w_{1}(l).\end{cases}$
It is worth noting that our ODE system (1.3) is similar to the one studied in [1]. However, the ODE system in [1] does not come from a discrete nonlinear wave equation.
Assumption 1.
We will assume that $M$ in (1.2) satisfies
$M>\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}.$ (1.4)
Theorem 1.1.
Assume (1.4). Let $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. There exists $u\in\mathcal{A}_{M}$ such that the solution to (1.3) satisfies,
$\begin{cases}x_{l}(t)=\bar{x}_{1l}+(t-T)\bar{v},\\ v_{l}(t)=\bar{v}=\sqrt{\alpha/\beta},\end{cases}$
for $l\in\overline{\Lambda}$, $t>T$, and some constant $\bar{x}_{1l}\in\mathbb{R}$. In particular, we have $w=\varphi_{1}$ for all $t\geq T$, where $\varphi_{1}$ is a moving flock with
$\varphi_{1}(l,t)=\bar{x}_{1l}+(t-T)\bar{v}=\bar{x}_{1l}-\bar{v}T+\bar{v}t\quad\text{ for }(l,t)\in\overline{\Lambda}\times\mathbb{R}.$
Theorem 1.2.
Assume (1.4). Let $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Let $\varphi_{2}(l,t)=\bar{\varphi}_{2}(l)+\bar{v}t$ be an arbitrary flock, where $\bar{\varphi}_{2}$ satisfies
$|-\Delta\bar{\varphi}_{2}(l)|<M\quad\text{ for }l\in\overline{\Lambda}$ (1.5)
Then, there exist $u\in\mathcal{A}_{M}$ and a constant $a\in\mathbb{R}$ such that the solution to (1.3) satisfies
$\begin{cases}x_{l}(t)=\bar{\varphi}_{2}(l)+a+\bar{v}t,\\ v_{l}(t)=\bar{v}=\sqrt{\alpha/\beta}\end{cases}$
for $l\in\overline{\Lambda}$ and $t>T$. In particular, we have $w=\varphi_{2}+a$ for all $t\geq T$.
As a by-product of Theorems 1.11.2, we have the following result.
Theorem 1.3.
Assume (1.4). Let $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Let $\varphi_{3}(l,t)=\bar{\varphi}_{3}(l)$ be an arbitrary stationary flock, where $\bar{\varphi}_{3}$ satisfies
$|-\Delta\bar{\varphi}_{3}(l)|<M\quad\text{ for }l\in\overline{\Lambda}.$ (1.6)
Then, there exist $u\in\mathcal{A}_{M}$ and a constant $a\in\mathbb{R}$ such that the solution to (1.3) satisfies
$\begin{cases}x_{l}(t)=\bar{\varphi}_{3}(l)+a,\\ v_{l}(t)=\bar{v}=0.\end{cases}$
for $l\in\overline{\Lambda}$ and $t>T$. In particular, we have $w=\varphi_{3}+a$ for all $t\geq T$.
Theorem 1.1 can, in some sense, be regarded as the “wave” counterpart to the “particle” results established in the seminal work [1] on controlling the Cucker-Smale model. In contrast, Theorems 1.2 and 1.3 are characteristic of wave phenomena and do not possess a direct “particle” analog. We note that the assumption (1.4) on $M$ in Theorems 1.11.3 is essentially optimal. See Remark 2.
Organization of the paper
The paper is organized as follows. In Section 2, we prove Theorem 1.1. The proofs of Theorems 1.21.3 are given in Section 3. Finally, section 4 contains a connection between the flocking problem and a mimimal time problem in the framework of optimal control theory and Hamilton–Jacobi equations.
2. Proof of Theorem 1.1
Our strategy to prove Theorem 1.1 is as follows. We start by using a control $u$ that drives the system toward zero velocity. To do so, we first design $u$ so that the particles’ velocities become sufficiently small after some finite time. Once the system is close to zero velocity, we can find a control in $\mathcal{A}_{M}$ that brings it to exactly zero velocity, a stationary flock, after another finite amount of time. Finally, we apply a third type of control to steer the system from a stationary flock to a moving flock with the desired group velocity $\bar{v}=\sqrt{\alpha/\beta}$. It is important noting that all the controls designed are explicit in terms of $(x_{l}(t))_{l\in\overline{\Lambda}}$ and $(v_{l}(t))_{l\in\overline{\Lambda}}$.
In the process of driving the system to zero velocity, we will study the Lyapunov function
$V(t)=\frac{1}{2}\sum_{l\in\overline{\Lambda}}\left[|v_{l}(t)|^{2}+\frac{1}{2}\sum_{l^{\prime}\sim l}\frac{(x_{l^{\prime}}(t)-x_{l}(t))^{2}}{h^{2}}\right].$
It is immediate that $V\geq 0$. We will now prove
$\dot{V}(t)=\sum_{l\in\overline{\Lambda}}\left[(\alpha-\beta|v_{l}(t)|^{2})|v_{l}(t)|^{2}+(u_{l}(t),v_{l}(t))\right]$
which shows that choosing $u_{l}(t)$ we can control $\dot{V}$. For clarity of our writing, we use the usual way to write dot product $(a,b)=ab$ for $a,b\in\mathbb{R}$.
We have
$\dot{V}(t)$ $=\sum_{l\in\overline{\Lambda}}\left[(\dot{v}_{l}(t),v_{l}(t))+\frac{1}{2}\sum_{l^{\prime}\sim l}\left(\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}},v_{l^{\prime}}(t)-v_{l}(t)\right)\right]$
$=\sum_{l\in\overline{\Lambda}}\Bigg{[}\left(\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}+(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)+u_{l}(t),v_{l}(t)\right)$
$\qquad\qquad\qquad+\frac{1}{2}\left(\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}},-v_{l}(t)\right)+\frac{1}{2}\sum_{l^{\prime}\sim l}\left(\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}},v_{l^{\prime}}(t)\right)\Bigg{]}$
$=\sum_{l\in\overline{\Lambda}}\left((\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)+u_{l}(t),v_{l}(t)\right)$
$=\sum_{l\in\overline{\Lambda}}\left[(\alpha-\beta|v_{l}(t)|^{2})|v_{l}(t)|^{2}+(u_{l}(t),v_{l}(t))\right].$
Let $\gamma>\max\left\{1,\frac{1}{M}\sqrt{\frac{\alpha^{3}}{\beta}}\right\}$ be fixed and define
$a_{1}=\frac{1}{\gamma}\sqrt{\frac{\alpha}{\beta}}\qquad\text{ and }\qquad a_{2}=\gamma\sqrt{\frac{\alpha}{\beta}}.$
When $t\in[0,T_{0}]$ for some $T_{0}$ we will define later, we will use the control $u_{l}$ based on $v_{l}$ in the following way:
$u_{l}(v_{l})=\begin{cases}0&\text{if }|v_{l}|\geq 2a_{2},\\ -M\frac{v_{l}}{|v_{l}|}\left(2-\frac{|v_{l}|}{a_{2}}\right)&\text{if }|v_{l}|\in(a_{2},2a_{2}),\\ -M\frac{v_{l}}{|v_{l}|}&\text{if }|v_{l}|\in[a_{1},a_{2}],\\ -M\frac{v_{l}}{a_{1}}&\text{if }|v_{l}|<a_{1}.\end{cases}$ (2.1)
Of course, $u_{l}$ is Lipschitz in $v_{l}$. For practical purposes, we do not need to turn on the control when $|v_{l}|\geq 2a_{2}$. See Remark 1. For each time $t$, we can split the indices $l$ into 3 sets:
$\begin{cases}I_{2}(t)=\{l\in\overline{\Lambda}\,:\,|v_{l}(t)|>a_{2}\},\\ I_{1}(t)=\{l\in\overline{\Lambda}\,:\,|v_{l}(t)|\in[a_{1},a_{2}]\},\\ I_{0}(t)=\{l\in\overline{\Lambda}\,:\,|v_{l}(t)|<a_{1}\}.\end{cases}$
For $l\in I_{2}(t)$, $u_{l}v_{l}=0$ if $|v_{l}|\geq 2a_{2}$ or
$u_{l}v_{l}=-M\frac{v_{l}}{|v_{l}|}\left(2-\frac{|v_{l}|}{a_{2}}\right)v_{l}=-M|v_{l}|\left(2-\frac{|v_{l}|}{a_{2}}\right)<0$
if $|v_{l}|\in(a_{2},2a_{2})$. As $\gamma>1$,
$(\alpha-\beta|v_{l}(t)|^{2})|v_{l}(t)|^{2}<(\alpha-\beta a_{2}^{2})|v_{l}(t)|^{2}<(\alpha-\gamma^{2}\alpha)|v_{l}(t)|^{2}<0.$
Hence,
$(\alpha-\beta|v_{l}(t)|^{2})|v_{l}(t)|^{2}+u_{l}v_{l}<(\alpha-\gamma^{2}\alpha)a_{2}^{2}.$
We define
$C_{2}:=(\gamma^{2}-1)\alpha a_{2}^{2}>0.$
For $l\in I_{1}(t)$, we see $u_{l}v_{l}=-M|v_{l}|<0$. Also,
$(\alpha-\beta|v_{l}|^{2})|v_{l}|^{2}+u_{l}v_{l}=|v_{l}|(\alpha|v_{l}|-\beta|v_{l}|^{3}-M).$ (2.2)
Consider the function $f(x)=\alpha x-\beta x^{3}$. It is clear that
$\max_{x\geq 0}f(x)=f\left(\sqrt{\frac{\alpha}{3\beta}}\right)=\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}<M$ (2.3)
by our assumption (1.4). Applying this to (2.2), we see
$(\alpha-\beta|v_{l}|^{2})|v_{l}|^{2}+u_{l}v_{l}<|v_{l}|\left(\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}-M\right)<a_{1}\left(\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}-M\right).$ (2.4)
We define
$C_{1}:=a_{1}\left(M-\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}\right)>0.$
Finally, for $l\in I_{0}(t)$, $u_{l}v_{l}=-M\frac{|v_{l}|^{2}}{a_{1}}$. Since $\gamma>\frac{1}{M}\sqrt{\frac{\alpha^{3}}{\beta}}$, we have $a_{1}<\frac{M}{\alpha}$. Thus,
$(\alpha-\beta|v_{l}|^{2})|v_{l}|^{2}+u_{l}v_{l}=|v_{l}|^{2}\left(\alpha-\beta|v_{l}|^{2}-\frac{M}{a_{1}}\right)\leq-\beta|v_{l}|^{4}.$
Therefore, we have shown that when $u_{l}$ is as described in (2.1),
$\dot{V}(t)\leq\sum_{l\in I_{2}(t)}-C_{2}+\sum_{l\in I_{1}(t)}-C_{1}+\sum_{l\in I_{0}(t)}-\beta|v_{l}(t)|^{4}.$ (2.5)
Thus, $V(t)$ is nonincreasing and in particular, $V(t)$ is bounded from above by $V(0)$. Since both terms in $V$ are nonnegative, this means they both must be bounded from above. Additionally, for $l\in\overline{\Lambda}$,
$\sum_{l^{\prime}\sim l}\frac{(x_{l^{\prime}}(t)-x_{l}(t))^{2}}{h^{2}}\leq V(0),$
which implies that
$\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}\right|\leq C.$ (2.6)
Therefore, we have
$|\dot{v}_{l}(t)|\leq C(1+\alpha+|u_{l}(t)|)\leq C\quad\text{or}\quad|\dot{v}_{l}(t)|\leq C(1+\beta C^{2}+|u_{l}(t)|)\leq C.$
Here, $C=C(V(0),h,\alpha,\beta)>0$.
Lemma 2.1.
The system (1.3) satisfies
$\lim_{t\to\infty}v_{l}(t)=0\quad\text{ for all $l\in\overline{\Lambda}$}$
when $u_{l}$ is given by (2.1).
Proof.
Assume this is not the case. Then there exists a $c_{0}>0$, an index $i$, and a sequence of times $t_{k}\to\infty$ such that
$|v_{l}(t_{k})|>c_{0}.$
Since $|\dot{v}_{l}(t)|\leq C$, there is a $\tau>0$ such that
$|v_{l}(t)|>\frac{c_{0}}{2}\quad\text{ for all $t\in(t_{k}-\tau,t_{k}+\tau)$, $k\in\mathbb{N}$.}$
Then
$\dot{V}(t)<\max\left\{-C_{2},-C_{1},-\beta\frac{c_{0}^{4}}{16}\right\}\quad\text{ for all $t\in(t_{k}-\tau,t_{k}+\tau)$, $k\in\mathbb{N}$.}$
This shows that
$\lim_{t\to\infty}V(t)=-\infty,$
which contradicts the fact that $V\geq 0$. Thus, $\lim_{t\to\infty}v_{l}(t)=0$. ∎
Remark 1.
In (2.1) above, we define $u_{l}=u_{l}(v_{l})$ in such a way that we do not need to use the control $u_{l}$ when $|v_{l}|\geq 2a_{2}$. This choice is not necessary, but in practical purposes having the control set to zero might be preferred and letting the system be driven by the large velocities in the system might be preferred. Another way, which is simpler but requires a higher magnitude of control, is to define $u_{l}$ as
$u_{l}(v_{l})=\begin{cases}-M\frac{v_{l}}{|v_{l}|}&\text{if }|v_{l}|\geq a_{1},\\ -M\frac{v_{l}}{a_{1}}&\text{if }|v_{l}|<a_{1}.\end{cases}$
The proof that this control drives the system towards zero velocity would be identical to the proof above except for the definition of $I_{1}(t)$.
Remark 2.
We note that (2.2) and (2.4) show the need for the assumption on $M$. The magnitude of $u$ has to be bounded by $M$. However, if we would like $\dot{V}<0$, then $u$ must have a large enough magnitude to cancel out $(\alpha-\beta|v_{l}|^{2})|v_{l}|$ which is equal to $\frac{2}{3\sqrt{2}}\sqrt{\frac{\alpha^{3}}{\beta}}$ when $|v_{l}|=\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}$. Thus, $M>\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}$ ensures $\dot{V}<0$. If $t\mapsto V(t)$ is not decreasing, Lemma 2.1 might not hold true.
Lemma 2.2.
The system (1.3) satisfies
$\lim_{t\to\infty}\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}=0\quad\text{ for all $l\in\overline{\Lambda}$}$ (2.7)
when $u_{l}$ is given by (2.1).
Proof.
Since $|\dot{x}_{l}(t)|=|v_{l}(t)|\leq\sqrt{2V(0)}$, the functions $x_{l}(t)$ are Lipschitz with constant $L:=\sqrt{2V(0)}$. Then $x_{l^{\prime}}(t)-x_{l}(t)$ are Lipschitz with Lipschitz constants less than or equal to $2L$. Assume now for the sake of contradiction that (2.7) does not hold. In light of (2.6), there exists an $l$ and a sequence $\{t_{k}\}$ such that $t_{k}\to\infty$ and
$\lim_{k\to\infty}\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t_{k})-x_{l}(t_{k})}{h^{2}}=c\neq 0.$ (2.8)
There exists a constant $K>0$ such that when $k>K$
$\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}-c\right|$
=\,= $\left|\left[\sum_{l^{\prime}\sim l}\frac{(x_{l^{\prime}}(t)-x_{l}(t))-((x_{l^{\prime}}(t_{k})-x_{l}(t_{k}))+(x_{l^{\prime}}(t_{k})-x_{l}(t_{k}))}{h^{2}}\right]-c\right|$
[…]
Additionally, there exists a sequence $(\eta_{k})_{k\geq K}$ such that when $t>t_{k}$,
$\begin{cases}|v_{l}(t)|<\eta_{k}<a_{1},\\ |(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)|<\eta_{k}<a_{1}\end{cases}$
and $\eta_{k}\to 0$ due to Lemma 2.1. Thus,
$\begin{split}&|\dot{v}_{l}(t_{k}+\tau)-c|\\ =\,&\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t_{k}+\tau)-x_{l}(t_{k}+\tau)}{h^{2}}-c+(\alpha-\beta|v_{l}(t_{k}+\tau)|^{2})v_{l}(t_{k}+\tau)+u_{l}(t_{k}+\tau)\right|\\ \leq\,&8nL\tau+\frac{|c|}{8}+\eta_{k}+\frac{M}{a_{1}}\eta_{k}.\end{split}$
Using the above inequality, we obtain
$|v_{l}(t_{k}+\tau)-c\tau|-|v_{l}(t_{k})|\leq|v_{l}(t_{k}+\tau)-v_{l}(t_{k})-c\tau|$
[…]
Thus,
$|v_{l}(t_{k}+\tau)-c\tau|$
[…]
Using the reverse triangle inequality again, we have
$|c\tau|-|v_{l}(t_{k}+\tau)|\leq\eta_{k}+8nL\tau^{2}+\frac{|c|\tau}{8}+\left(1+\frac{M}{a_{1}}\right)\tau\eta_{k}.$
Now if $\tau=\min\Big{\{}\frac{|c|}{16nL},1\Big{\}}$, then $8nL\tau^{2}\leq\frac{|c|}{2}\tau$, so
$\frac{3|c|\tau}{8}-\tau\eta_{k}\left(1+\frac{M}{a_{1}}\right)-\eta_{k}\leq|v_{l}(t_{k}+\tau)|\quad\text{ for all }k>K.$ (2.9)
Since $\eta_{k}\to 0$, we know there exists $K_{2}$ such that when $k>K_{2}$, $\eta_{k}(1+\frac{M}{a_{1}})\leq\frac{|c|}{16}$ and $\eta_{k}<\frac{|c|\tau}{16}$. Thus, when $k>\max\{K,K_{2}\}$,
$|v_{l}(t_{k}+\tau)|\geq\frac{|c|\tau}{4}.$
However, if we take the limit as $k\to\infty$, we have $\frac{|c|\tau}{4}\leq 0$. Therefore, $c=0$, which is contradicts (2.8). The proof is complete. ∎
Therefore, there exists a $T_{0}$ such that if (2.1) is used for $0<t\leq T_{0}$,
$|v_{l}(T_{0})|<\epsilon\quad\text{ and }\quad\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{0})-x_{l}(T_{0})}{h^{2}}\right|<\epsilon$
for some $\epsilon$ satisfying
$0<\epsilon<\min\left\{1,\frac{M}{2n+1+\frac{8n}{h^{2}}+2\alpha+8\beta}\right\}.$ (2.10)
It is important noting that $\epsilon$ can be chosen explicit here with the above bound, which is useful in practice. We are now ready to finish the proof of Theorem 1.1.
Proof of Theorem 1.1.
Define $T_{1,l}=T_{0}+\frac{1}{\epsilon}|v_{l}(T_{0})|$ for each $l\in\overline{\Lambda}$ and $T_{1}=\max_{l\in\overline{\Lambda}}T_{1,l}$. For $T_{0}<t\leq T_{1}$, we define
$u_{l}(t)=\dot{v}_{l}(t)-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}-(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)$ (2.11)
where $v_{l}(t)$ satisfies the following:
$v_{l}(t)=\begin{cases}v_{l}(T_{0})-\epsilon(t-T_{0})\frac{v_{l}(T_{0})}{|v_{l}(T_{0})|}\quad&\text{for }t\in[T_{0},T_{1,l}],\\ 0&\text{for }t>T_{1,l}.\end{cases}$ (2.12)
and
[…]
$=x_{l}(T_{0})+v_{l}(T_{0})(t-T_{0})-\frac{\epsilon}{2}\frac{v_{l}(T_{0})}{|v_{l}(T_{0})|}(t-T_{0})^{2}$
when $t\in[T_{0},T_{1,l}]$ and $x_{l}(t)=x_{l}(T_{1,l})$ when $t>T_{1,l}$. Notice
$|x_{l}(t)-x_{l}(T_{0})|$ $\leq|v_{l}(T_{0})|(T_{1,l}-T_{0})+\frac{\epsilon}{2}(T_{1,l}-T_{0})^{2}$
$\leq\frac{|v_{l}(T_{0})|^{2}}{\epsilon}+\frac{\epsilon}{2}\frac{|v_{l}(T_{0})|^{2}}{\epsilon^{2}}<2\epsilon.$
We must show that $|u_{l}|<M$. We first observe that when $t\in(T_{1,l},T_{1}]$,
$u_{l}(t)=0-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1,l})-x_{l}(T_{1,l})}{h^{2}}-0=-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1,l})-x_{l}(T_{1,l})}{h^{2}}.$
Thus,
$|u_{l}(t)|$ $=\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1,l})-x_{l}(T_{1,l})}{h^{2}}\right|$
$=\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1,l})-x_{l^{\prime}}(T_{0})+x_{l^{\prime}}(T_{0})-x_{l}(T_{0})+x_{l}(T_{0})-x_{l}(T_{1,l})}{h^{2}}\right|$
$\leq\sum_{l^{\prime}\sim l}\left(\frac{2\epsilon}{h^{2}}+\epsilon+\frac{2\epsilon}{h^{2}}\right)=\frac{8n\epsilon}{h^{2}}+2n\epsilon$
when $t\in(T_{1,l},T_{1}].$ Similarly, when $t\in(T_{0},T_{1,l}]$,
$u_{l}(t)$ $=-\epsilon\frac{v_{l}(T_{0})}{|v_{l}(T_{0})|}-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}$
$\qquad-\left(\alpha-\beta\left|v_{l}(T_{0})-\epsilon(t-T_{0})\frac{v_{l}(T_{0})}{|v_{l}(T_{0})|}\right|^{2}\right)\left(v_{l}(T_{0})-\epsilon(t-T_{0})\frac{v_{l}(T_{0})}{|v_{l}(T_{0})|}\right).$
Hence, using the bound on $\epsilon$ given in (2.10) we see
$|u_{l}(t)|$
[…]
=\,= $\epsilon\left(2n+1+\frac{8n}{h^{2}}+2\alpha+8\beta\epsilon^{2}\right)<M.$
Thus, we have designed a control that drives the system to zero velocity at $T_{1}$. The system is now at a stationary flock at $t=T_{1}$.
Finally, let
$u_{l}(t)=-(\alpha-\beta|v_{l}|^{2})v_{l}-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}+w\quad\text{ for $t\in[T_{1},T_{2}]$,}$
where
$T_{2}=T_{1}+\frac{2\sqrt{\alpha}}{M\sqrt{\beta}}\quad\text{ and }\quad w=\frac{\bar{v}}{T_{2}-T_{1}}=\frac{M}{2}.$
We notice that the first terms of $u_{l}$ will cancel out the other terms on the right side of the second equation in (1.3) so that
$\dot{v}_{l}(t)=w=\frac{M}{2}.$
Additionally, $v_{l}(T_{1})=0$. Thus, $v_{l}(t)$ is the same for all $l\in\overline{\Lambda}$ and
$v_{l}(t)=\frac{M(t-T_{1})}{2}\quad\text{ for }t\in[T_{1},T_{2}].$
Therefore,
$\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}=\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1})-x_{l}(T_{1})}{h^{2}}\quad\text{ for all $t\in[T_{1},T_{2}]$.}$
Recall that this term has magnitude less than $\epsilon\ll M/2$. With this information, we can show that $|u_{l}|<M$. As $0\leq v_{l}(t)\leq\bar{v}$ we can see using (2.3) that
$0\leq(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)\leq\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}<M.$
Thus,
$|u_{l}(t)|\leq\left|-(\alpha-\beta|v_{l}|^{2})v_{l}+\frac{M}{2}\right|+\left|-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}\right|\leq\frac{M}{2}+\epsilon<M.$
At $t=T_{2}$, we get
$v_{l}(T_{2})=\bar{v}=\sqrt{\frac{\alpha}{\beta}},$
and
$\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{2})-x_{l}(T_{2})}{h^{2}}\right|=\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1})-x_{l}(T_{1})}{h^{2}}\right|\leq\epsilon.$
For $t\geq T_{2}$, we simply keep $v_{l}(t)=v_{l}(T_{2})=\bar{v}$ for all $l\in\overline{\Lambda}$. Then, we get
$x_{l}(t)=x_{l}(T_{2})+\bar{v}(t-T_{2})\quad\text{ for all }t\geq T_{2}.$
The proof is complete with $T=T_{2}$ and $\bar{x}_{1l}=x_{l}(T_{2})$ for $l\in\overline{\Lambda}$. ∎
Remark 3.
It is important to note that, in the above proof, for $l\in\overline{\Lambda}$,
$\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{2})-x_{l}(T_{2})}{h^{2}}\right|=\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1})-x_{l}(T_{1})}{h^{2}}\right|\leq\epsilon,$
where $\epsilon>0$ can be chosen as small as we want. This is an important point that we will need to use in the proof of Theorem 1.2.
3. Proofs of Theorems 1.21.3
3.1. Proof of Theorem 1.2
Our strategy to prove Theorem 1.2 is as follows. By Theorem 1.1, we have driven the system to a moving flock at time $T_{2}$. By designing an explicit control, we drive the system to a moving flock where every particle is at the same position at $T_{3}>T_{2}$. From this intermediate moving flock where all particles are at the same position, we will finally design another explicit control driving the system to the desired flock at $T_{4}>T_{3}$. Along the way, we need to make sure that the changes of the positions of the particles and their corresponding velocities are sufficiently small so that the controls always satisfy the given bound $M$.
Proof of Theorem 1.2.
By the proof of Theorem 1.1, we know that given any initial conditions, we can drive the system to a moving flock $\varphi_{1}(l,t)=\bar{\varphi}_{1}(l)+\bar{v}t$ at time $t=T_{2}$ satisfying
$\left|\sum_{l^{\prime}\sim l}\frac{\bar{\varphi}_{1}(l^{\prime})-\bar{\varphi}_{1}(l)}{h^{2}}\right|\leq\epsilon$ (3.1)
for all $l\in\overline{\Lambda}$. We will then adjust the control to drive the system to
$\begin{cases}x_{l}(T_{3})=x_{j}(T_{3}),\\ v_{l}(T_{3})=v_{j}(T_{3})=\bar{v},\end{cases}$
for all $l,j\in\overline{\Lambda}$. That is, the system is at a moving flock where every particle is at the same position at $T_{3}>T_{2}$. Let
$\bar{y}=\frac{1}{D^{n}}\sum_{l\in\overline{\Lambda}}\bar{\varphi}_{1}(l),$
and let $T_{3}>T_{2}$ satisfy
$\begin{split}&\frac{3\alpha|\bar{y}-\bar{\varphi}_{1}(l)|(T_{3}-T_{2})^{2}+6.75\sqrt{\alpha\beta}|\bar{y}-\bar{\varphi}_{1}(l)|^{2}(T_{3}-T_{2})}{(T_{3}-T_{2})^{3}}\\ &+\frac{3.375\beta|\bar{y}-\bar{\varphi}_{1}(l)|^{3}+6|\bar{y}-\bar{\varphi}_{1}(l)|(T_{3}+T_{2})}{(T_{3}-T_{2})^{3}}<M-\epsilon\end{split}$ (3.2)
for all $l\in\overline{\Lambda}$.
Remark 4.
We note that our choice of $\bar{y}$ is reasonable, but another point could be chosen. In fact, if we choose $\bar{y}$ such that it minimizes $\max_{l\in\Lambda}|y-\bar{\varphi}_{1}(l)|$ then we can reduce $T_{3}$.
We then set
$\begin{split}u_{l}(t)=&-\left[1+\frac{2(t-T_{2})^{3}}{(T_{3}-T_{2})^{3}}-\frac{3(t-T_{2})^{2}}{(T_{3}-T_{2})^{2}}\right]\sum_{l^{\prime}\sim l}\frac{\bar{\varphi_{1}}(l^{\prime})-\bar{\varphi_{1}}(l)}{h^{2}}\\ &+\frac{12\alpha}{(T_{3}-T_{2})^{3}}(\bar{y}-\bar{\varphi}_{1}(l))(t-T_{2})(T_{3}-t)\\ &+\frac{108\sqrt{\alpha\beta}}{(T_{3}-T_{2})^{6}}(\bar{y}-\bar{\varphi}_{1}(l))^{2}(t-T_{2})^{2}(T_{3}-t)^{2}\\ &+\frac{216\beta}{(T_{3}-T_{2})^{9}}(\bar{y}-\bar{\varphi}_{1}(l))^{3}(t-T_{2})^{3}(T_{3}-t)^{3}+\frac{6(\bar{y}-\bar{\varphi_{1}}(l))}{(T_{3}-T_{2})^{3}}(T_{3}+T_{2}-2t)\end{split}$
for $t\in[T_{2},T_{3}]$. With this choice of control and the fact that $v(T_{2})=\bar{v}$, we find
$v_{l}(t)=\bar{v}+\frac{6(\bar{y}-\bar{\varphi_{1}}(l))}{(T_{3}-T_{2})^{3}}(t-T_{2})(T_{3}-t)$
and
$x_{l}(t)=\bar{\varphi}_{1}(l)+(t-T_{2})\bar{v}+\frac{3(t-T_{2})^{2}}{(T_{3}-T_{2})^{2}}(\bar{y}-\bar{\varphi_{1}}(l))-\frac{2(t-T_{2})^{3}}{(T_{3}-T_{2})^{3}}(\bar{y}-\bar{\varphi_{1}}(l)).$
Since $x_{l}(T_{3})=\bar{y}+(T_{3}-T_{2})\bar{v}$, we deduce $x_{l}(T_{3})=x_{j}(T_{3})$ and $v_{l}(T_{3})=v_{j}(T_{3})=\bar{v}$ for all $l,j\in\overline{\Lambda}$. However, we must show that $|u_{l}(t)|<M$. Notice that $(t-T_{2})(T_{3}-t)$ has a maximum value of $\frac{(T_{3}-T_{2})^{2}}{4}$ and
$1+\frac{2(t-T_{2})^{3}}{(T_{3}-T_{2})^{3}}-\frac{3(t-T_{2})^{2}}{(T_{3}-T_{2})^{2}}\leq 1\quad\text{ for $t\in[T_{2},T_{3}]$.}$
Combining these facts with (3.6) and (LABEL:T2cond) shows
$|u_{l}(t)|$
[…]
$\,+\frac{108\sqrt{\alpha\beta}}{(T_{3}-T_{2})^{6}}|\bar{y}-\bar{\varphi}_{1}(l)|^{2}\frac{(T_{3}-T_{2})^{4}}{16}+\frac{216\beta}{(T_{3}-T_{2})^{9}}|\bar{y}-\bar{\varphi}_{1}(l)|^{3}\frac{(T_{3}-T_{2})^{6}}{64}$
$\,+\frac{6(\bar{y}-\bar{\varphi}_{1}(l))}{(T_{3}-T_{2})^{3}}(T_{3}+T_{2})$
[…]
$\,\,+\frac{3.375\beta|\bar{y}-\bar{\varphi}_{1}(l)|^{3}+6|\bar{y}-\bar{\varphi}_{1}(l)|(T_{3}+T_{2})}{(T_{3}-T_{2})^{3}}<M.$
From the intermediate moving flock where all particles are at the same position, we will finally design the control for $t\in[T_{3},T_{4}]$ that drives the system to the desired flock
$x_{l}(t)=\bar{\varphi}_{2}(l)+a+\bar{v}t\quad\text{ for }l\in\overline{\Lambda},t\geq T_{4}.$
By the assumptions, $|-\Delta\bar{\varphi}_{2}|<M$. Let $T_{4}>T_{3}$ satisfy
$\begin{split}&\frac{3\alpha|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|(T_{4}-T_{3})^{2}+6.75\sqrt{\alpha\beta}|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|^{2}(T_{4}-T_{3})}{(T_{4}-T_{3})^{3}}\\ &+\frac{3.375\beta|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|^{3}+6|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|(T_{4}+T_{3})}{(T_{4}-T_{3})^{3}}<M-|\Delta\bar{\varphi}_{2}(l)|\end{split}$ (3.3)
for all $l\in\overline{\Lambda}$.
Now from $T_{3}$ to $T_{4}$, letting
$\begin{split}u_{l}(t)=&\left[-\frac{3(t-T_{3})^{2}}{(T_{4}-T_{3})^{2}}+\frac{2(t-T_{3})^{3}}{(T_{4}-T_{3})^{3}}\right]\sum_{l^{\prime}\sim l}\frac{\bar{\varphi}_{2}(l^{\prime})-\bar{\varphi}_{2}(l)}{h^{2}}\\ &+\frac{12\alpha}{(T_{4}-T_{3})^{3}}(\bar{\varphi}_{2}(l)-x_{l}(T_{3}))(t-T_{3})(T_{4}-t)\\ &+\frac{108\sqrt{\alpha\beta}}{(T_{4}-T_{3})^{6}}(\bar{\varphi}_{2}(l)-x_{l}(T_{3}))^{2}(t-T_{3})^{2}(T_{4}-t)^{2}\\ &+\frac{216\beta}{(T_{4}-T_{3})^{9}}(\bar{\varphi}_{2}(l)-x_{l}(T_{3}))^{3}(t-T_{3})^{3}(T_{4}-t)^{3}\\ &+\frac{6(\bar{\varphi}_{2}(l)-x_{l}(T_{3}))}{(T_{4}-T_{3})^{3}}(T_{4}+T_{3}-2t),\end{split}$
we arrive at
$v_{l}(t)=\bar{v}+\frac{6(\bar{\varphi}_{2}(l)-x_{l}(T_{3}))}{(T_{4}-T_{3})^{3}}(t-T_{3})(T_{4}-t)$
and
$x_{l}(t)=x_{l}(T_{3})+(t-T_{3})\bar{v}+\frac{3(t-T_{3})^{2}}{(T_{4}-T_{3})^{2}}(\bar{\varphi}_{2}(l)-x_{l}(T_{3}))-\frac{2(t-T_{3})^{3}}{(T_{4}-T_{3})^{3}}(\bar{\varphi}_{2}(l)-x_{l}(T_{3})).$
Then,
$\begin{cases}x_{l}(T_{4})=(T_{4}-T_{3})\bar{v}+\bar{\varphi}_{2}(l),\\ v_{l}(T_{4})=\bar{v}.\end{cases}$
Once again, we would like to ensure $|u_{l}(t)|<M$. For $l\in\overline{\Lambda}$, we estimate
$|u_{l}(t)|$
[…]
$+\frac{3.375\beta|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|^{3}+6|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|(T_{4}+T_{3})}{(T_{4}-T_{3})^{3}}<M.$
Finally, for $t>T_{4}$, we will let
$u_{l}(t)=-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}},$
so $v_{l}(t)=\bar{v}$, and $x_{l}(t)=(t-T_{3})\bar{v}+\bar{\varphi}_{2}(l)$ for $t\geq T_{4}$. Here, $a=-T_{3}\bar{v}$. ∎
Remark 5.
We note that assumption (1.5) in Theorem 1.2 can be relaxed to either
$-M\leq-\Delta\bar{\varphi}_{2}(l)<M\quad\text{ for all }l\in\overline{\Lambda}$ (3.4)
or
$-M<-\Delta\bar{\varphi}_{2}(l)\leq M\quad\text{ for all }l\in\overline{\Lambda}.$ (3.5)
Indeed, let us assume (3.4). In this case, we take
$\bar{z}=-1+\min_{l\in\overline{\Lambda}}\left(x_{l}(T_{3})-\bar{\varphi}_{2}(l)\right).$
Let $T_{4}>T_{3}$ satisfy
$\begin{split}&\frac{3\alpha|\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3})|(T_{4}-T_{3})^{2}+6.75\sqrt{\alpha\beta}|\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3})|^{2}(T_{4}-T_{3})}{(T_{4}-T_{3})^{3}}\\ &+\frac{3.375\beta|\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3})|^{3}+6|\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3})|(T_{4}+T_{3})}{(T_{4}-T_{3})^{3}}<M+\Delta\bar{\varphi}_{2}(l)\end{split}$
for all $l\in\overline{\Lambda}$. Denoting by
$v_{l}(t)=\bar{v}+\frac{6(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))}{(T_{4}-T_{3})^{3}}(t-T_{3})(T_{4}-t)\leq\bar{v}\quad\text{ for }t\in[T_{3},T_{4}],$
we obtain
$\dot{v}_{l}(t)=\frac{6(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))}{(T_{4}-T_{3})^{3}}(T_{3}+T_{4}-2t)\geq 0\quad\text{ for }t\in[(T_{3}+T_{4})/2,T_{4}].$
Then, for $t\in[(T_{3}+T_{4})/2,T_{4}]$ which is close to $T_{4}$,
$\begin{split}u_{l}(t)=&\left[-\frac{3(t-T_{3})^{2}}{(T_{4}-T_{3})^{2}}+\frac{2(t-T_{3})^{3}}{(T_{4}-T_{3})^{3}}\right]\Delta\bar{\varphi}_{2}(l)\\ &+\frac{12\alpha}{(T_{4}-T_{3})^{3}}(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))(t-T_{3})(T_{4}-t)\\ &+\frac{108\sqrt{\alpha\beta}}{(T_{4}-T_{3})^{6}}(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))^{2}(t-T_{3})^{2}(T_{4}-t)^{2}\\ &+\frac{216\beta}{(T_{4}-T_{3})^{9}}(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))^{3}(t-T_{3})^{3}(T_{4}-t)^{3}\\ &+\frac{6(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))}{(T_{4}-T_{3})^{3}}(T_{4}+T_{3}-2t)\\ \geq&\left[-\frac{3(t-T_{3})^{2}}{(T_{4}-T_{3})^{2}}+\frac{2(t-T_{3})^{3}}{(T_{4}-T_{3})^{3}}\right]\Delta\bar{\varphi}_{2}(l)\\ &+\frac{6(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))}{(T_{4}-T_{3})^{3}}\left[(T_{4}+T_{3}-2t)+2\alpha(t-T_{3})(T_{4}-t)\right]\\ &+\frac{216\beta}{(T_{4}-T_{3})^{9}}(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))^{3}(t-T_{3})^{3}(T_{4}-t)^{3}\\ \geq&\left[-\frac{3(t-T_{3})^{2}}{(T_{4}-T_{3})^{2}}+\frac{2(t-T_{3})^{3}}{(T_{4}-T_{3})^{3}}\right]\Delta\bar{\varphi}_{2}(l)\geq-M.\\ \end{split}$
Therefore, we can conclude that
$-M\leq u_{l}(t)<M\quad\text{ for all }t\in[T_{3},T_{4}].$
Thus, Theorem 1.2 still holds if we replace (1.5) by either (3.4) or (3.5).
3.2. Proof of Theorem 1.3
Sketch of the proof of Theorem 1.3.
Since this is a by-product of Theorems 1.11.2, we only give a sketch of the proof here.
By the proof of Theorem 1.1, we know that given any initial conditions, we can drive the system to a stationary flock at time $t=T_{1}$ satisfying
$\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1})-x_{l}(T_{1})}{h^{2}}\right|\leq\epsilon$ (3.6)
for all $l\in\overline{\Lambda}$.
We then use the first step in the proof of Theorem 1.2 to drive the system to a stationary flock where all particles are at the same position at time $T_{5}>T_{1}$. Afterwards, we use the second step in the proof of Theorem 1.2 to drive the system to the desired stationary flock at $T_{6}>T_{5}$. That is, for $l\in\overline{\Lambda}$ and $t\geq T_{6}$,
$\begin{cases}x_{l}(t)=\bar{\varphi}_{3}(l)+a,\\ v_{l}(t)=\bar{v}=0.\end{cases}$
4. Connection to optimal control and Hamilton–Jacobi equations
The ODE system (1.3) is clearly nonlocal in nature. It can be formulated in the optimal control framework as follows.
Write $X_{1}=(x_{l})_{l\in\overline{\Lambda}}$ and $X_{2}=(v_{l})_{l\in\overline{\Lambda}}$ as vectors in $\mathbb{R}^{D^{n}}$. Denote by $X=(X_{1},X_{2})\in\mathbb{R}^{2D^{n}}$. Then, (1.3) can be written as
$\dot{X}=\tilde{b}(X,W)=b(X)+(0,W),$
with given initial condition
$X(0)=\left((w_{0}(l))_{l\in\overline{\Lambda}},(w_{1}(l))_{l\in\overline{\Lambda}}\right).$
Here, the vector field $b$ is
$b(X)=\left(X_{2},\left(\sum_{l^{\prime}\sim l}\frac{(X_{1})_{l^{\prime}}-(X_{1})_{l}}{h^{2}}+(\alpha-\beta|(X_{2})_{l}|^{2})(X_{2})_{l}\right)_{l\in\overline{\Lambda}}\right).$
In this framework, $W$ is the control and the admissible set is
$\overline{\mathcal{A}}_{M}=\left\{W\in\mathbb{R}^{D^{n}}\,:\,|W_{l}|=|u_{l}|\leq M\qquad\text{ for }l\in\overline{\Lambda}\right\}.$
Basically, we can think of $X$ as the master state since it contains all information, and the ODE for $X$ is then local in nature.
Without loss of generality, we are only concerned with moving flocks here. For a given state $X$, we are interested in the optimal time it takes to reach a moving flock, that is,
$\Gamma=\left\{X\;:\,\left|\sum_{l^{\prime}\sim l}\frac{(X_{1})_{l^{\prime}}-(X_{1})_{l}}{h^{2}}\right|\leq M,(X_{2})_{l}=-\sqrt{\alpha/\beta}\quad\text{ for}l\in\overline{\Lambda}\right\}\\ \bigcup\left\{X\;:\,\left|\sum_{l^{\prime}\sim l}\frac{(X_{1})_{l^{\prime}}-(X_{1})_{l}}{h^{2}}\right|\leq M,(X_{2})_{l}=\sqrt{\alpha/\beta}\quad\text{ for }l\in\overline{\Lambda}\right\}.$
Let $U(X)$ be the minimal time it takes to start from state $X$ to reach $\Gamma$. Then, for $X=(X_{1},X_{2})\in\mathbb{R}^{2D^{n}}$ and $P=(P_{1},P_{2})\in\mathbb{R}^{2D^{n}}$, the corresponding Hamiltonian is
$H(X,P)=\sup_{W\in\overline{\mathcal{A}}_{M}}\left(-\tilde{b}(X,W)\cdot P-1\right)=-b(X)\cdot P+M|P_{2}|_{1}-1.$
Here, $|\cdot|_{1}$ is the usual $1$-norm in $\mathbb{R}^{D^{n}}$. Therefore, $U$ satisfies the following eikonal type equation in the viscosity sense
$\begin{cases}-b(X)\cdot D_{X}U(X)+M|D_{X_{2}}U(X)|_{1}=1\qquad&\text{ for }X\in\mathbb{R}^{2D^{n}},\\ U(X)=0\qquad&\text{ for }X\in\Gamma.\end{cases}$ (4.1)
See [13, Chapter 2]. We note that this is a very complicated PDE posed in high dimensions as the vector field $b$ contains much information about the dynamics. The value function $U(X)$ gives us the minimal time to reach a moving flock starting from $X\in\mathbb{R}^{2D^{n}}$, but does not give us directly an optimal strategy to reach a moving flock. By Theorem 1.1, we have that $U$ is finite. As the Hamiltonian $H$ here is noncoercive in $P$, the regularity of the viscosity solution $U$ is not yet known.
The advantage of Theorems 1.11.2 is that they provide more or less explicit strategies to reach a moving flock in finite time. Nevertheless, it is important to emphasize that an optimal control framework is hidden behind this problem.
Equation (4.1) is interesting in its own right and merits further study. Specifically, the regularity of the viscosity solution $U$, the properties of the optimal trajectories (strategies) and possible approximations are of particular interest.