### v3

# Controlling Klein-Gordon chains and lattices

Abstract.
In this work, we initiate the study of controlling nonlinear Klein–Gordon chains and lattices through their emergent collective flocking behavior. By constructing appropriate feedback control mechanisms, we demonstrate that any physically admissible flock state can be achieved in finite time, meaning the chain can be driven from arbitrary initial vibrations toward a coherent traveling-wave motion. Finally, we reveal a deep connection between the flocking problem and a minimal-time control principle formulated within the framework of nonlinear Hamilton–Jacobi equations and optimal control theory, providing a unifying viewpoint for wave control in discrete nonlinear media.
Key words and phrases:
discrete semilinear wave equations; flocks; controls; optimal control; Hamilton–Jacobi equations
2010 Mathematics Subject Classification:
35B65, 35F21, 35Q99, 44A10, 45J05
The work of $S$ is partially supported by NSF grant DMS-2037851. The work of HVT is partially supported by NSF grant DMS-2348305. The work of MBT is partially supported by NSF Grants CAREER DMS-2303146, DMS-2204795, DMS-2306379.
1. Introduction
1.1. Brief overview
Fermi, Pasta, Ulam, and Tsingou [18] investigated how systems evolve toward thermal equilibrium by studying a model consisting of a chain of particles coupled through anharmonic interactions. Their numerical experiment, reported in their original study, produced highly non-intuitive results and eventually led to what is now known as the Fermi–Pasta–Ulam–Tsingou (FPUT) paradox. Although the problem turned out to be far more subtle than they initially expected, their work generated several influential by-products, including the discovery of solitons and breathers, and the realization that FPUT dynamics can, in certain regimes, be approximated by integrable systems. For modern perspectives on these developments, see [1, 2]. The FPUT model is one of the earliest and most renowned examples of an anharmonic lattice. Thus, the FPUT model stands as a classic and foundational system in the study of anharmonic chains and lattices. Progress on the non-equilibrium statistical mechanics of anharmonic chains and lattices remained slow until the late 1990s, when numerical experiments revealed unmistakable signs of anomalous energy transport [3]. Some elements of this behavior had been observed earlier, but only after these findings did extensive, systematic studies begin across many research groups, culminating in the comprehensive reviews [4, 5, 2].
Control theory, developed extensively over the past six decades, has produced deep insights and powerful tools across a wide range of disciplines, including finance, medicine, mechanics, engineering, and climatology [33]. Atomic vibrations in anharmonic chains (one-dimensional systems) and lattices (multi-dimensional systems) are described by discrete wave-type equations [18, 22, 25, 26]. The control of anharmonic chains has emerged as an important research frontier, driving progress in both fundamental science and technological applications (see, for example, [6, 7, 8, 9, 10, 12, 13, 28, 29, 30, 31] and the references therein). This line of research is motivated by the fact that controlling such systems enables the manipulation, redirection, and stabilization of energy flow in discrete media, which are central challenges in condensed matter physics and materials science. Moreover, anharmonic chains are highly nonlinear and often exhibit complex behaviors such as chaos, metastability, and long-time fluctuations. Control mechanisms provide a means to prevent instabilities, suppress unwanted oscillations, stabilize special configurations (including flock states, breathers, and traveling waves), and ensure desirable long-time dynamics. These capabilities are essential for the design of stable devices involving micro- and nanoscale vibrations. Modern engineered systems, including photonic crystals, mechanical metamaterials, and micromechanical resonator arrays, often mirror the structure of anharmonic chains. As a result, control-theoretic tools play a key role in shaping wave propagation in metamaterials, designing energy-harvesting systems, creating tunable filters and waveguides, and managing thermal or mechanical stresses at small scales. Control theory for anharmonic lattices and chains continues to directly influence the development of emerging technologies.
Despite the aforementioned importance in science and technology, rigorous mathematical control designs for anharmonic chains and lattices through the associated discrete wave-type equations have appeared only recently, in [19, 20], where impulsive, feedback, and internal control strategies for harmonic chains were established. Continuing recent studies on the control of anharmonic lattices [19, 20], which focus on linear anharmonic lattices, our work aims to initiate a systematic investigation of how atomic vibrations in nonlinear anharmonic chains and lattices can be controlled through the corresponding nonlinear discrete Klein–Gordon wave equations (Klein–Gordon chains and lattices), following the formulation in [24] (see also [2, 23]).
Because wave vibrations on lattices often display intricate and nonlinear dynamics [11, 27], we focus on designing control mechanisms capable of stabilizing these physically relevant fixed-grid systems by steering them toward flock states, coherent traveling-wave configurations in which particles move collectively with a common velocity direction (see Definition 1). Our work provides the first mathematical foundations for rigorous control designs for nonlinear anharmonic chains and lattices, an important research direction that, despite its significance as discussed above, remains largely unexplored.
We remark that controlling discrete nonlinear wave equations for numerical purposes has been well studied in the literature [17, 34]. However, the goal of this research was to design a control that is a good approximation for the continuous problem. Thus, the goal is to observe the behavior of the control as the mesh size tends to zero. Since our problem is meant to be applied to physical lattices that have a fixed size, there are large differences in the theory.
1.2. Settings and main goals
Let $D,n\in\mathbb{N}$ be given natural numbers, and let $h=1/D$ be the fixed grid size. Denote by
$\Lambda=\left\{\frac{j}{D}\,:\,j\in\mathbb{Z}\right\}^{n}=\left\{jh\,:\,j\in\mathbb{Z}\right\}^{n}$
the discrete grid in $\mathbb{R}^{n}$.
We consider the following discrete semilinear Klein–Gordon equation with a cubic nonlinearity following the formulation in [24, Section 2, Equation (2.11) and Section 8, Equation (8.3)] and a control term
$\begin{cases}w_{tt}-\Delta_{D}w=(\alpha-\beta|w_{t}|^{2})w_{t}+u(l,t)\qquad&\text{ in }\Lambda\times(0,\infty),\\ w(\cdot,0)=w_{0},~w_{t}(\cdot,0)=w_{1}\qquad&\text{ on }\Lambda.\end{cases}$ (1.1)
Note that the semilinear Klein–Gordon equation with both cubic and quadratic nonlinearities can be used to describe atomic vibrations in anharmonic lattices [24]. In this work, we focus solely on the Klein–Gordon equation with a cubic nonlinearity, while the other case can be treated similarly.
Here, $\alpha,\beta>0$ are given constants, and $w_{0},w_{1}$ are the given initial data, which are $\mathbb{Z}^{n}$-periodic. Let $e_{k}$ for $k\in\{1,\ldots,n\}$ be the standard basis vectors. For $l\in\Lambda$, we write $l^{\prime}\sim l$ if $l^{\prime}-l=\pm he_{k}$ for some $k\in\{1,\ldots,n\}$. The corresponding discrete Laplacian is
$\Delta_{D}w(l,t)=\Delta w(l,t)=\sum_{l^{\prime}\sim l}\frac{w(l^{\prime},t)-w(l,t)}{h^{2}}.$
When there is no confusion, we write $\Delta_{D}w(l,t)=\Delta w(l,t)$ for simplicity. The function $u:\Lambda\times[0,\infty)\to\mathbb{R}$, which is $\mathbb{Z}^{n}$-periodic in $l$, is our control, and we typically assume that
$u\in\mathcal{A}_{M}=\{f\in L^{\infty}(\Lambda\times[0,\infty))\,:\,f\text{ is $\mathbb{Z}^{n}$-periodic in $l$, }\|f\|_{L^{\infty}}\leq M\}$
where $M>0$ is a given positive constant (to be specified later).
As we are in the spatial $\mathbb{Z}^{n}$-periodic setting, we are concerned with the spatial $\mathbb{Z}^{n}$-periodic solution $w$ to (1.1). In particular, our problem can be phrased in the discrete torus
$\overline{\Lambda}=\left\{0,\frac{1}{D},\ldots,\frac{D-1}{D}\right\}^{n},$
where we identify $k+j/D\equiv j/D$ for $k,j\in\mathbb{Z}$. If a function $f:\Lambda\to\mathbb{R}$ is $\mathbb{Z}^{n}$-periodic, we can think of $f$ as a function from $\overline{\Lambda}$ to $\mathbb{R}$ as well, and vice versa. As explained in the introduction, our mesh size $h$ is fixed, which is the same as the lattice considered by Peierls [22].
In this paper, we are interested in using the control $u$ to drive the solution $w$ of (1.1) to a flock where the group velocity is constant. Looking at (1.1), it is clear that the nonlinear term vanishes if
$w_{t}=\bar{v}=\pm\sqrt{\frac{\alpha}{\beta}}\qquad\text{ or }\qquad w_{t}=\bar{v}=0.$
Further, by using the ansatz $w(l,t)=\bar{w}(l)+\bar{v}t$ in (1.1), we see that
$|-\Delta\bar{w}(l)|=|u(l,t)|\leq M\quad\text{ for }l\in\overline{\Lambda},$ (1.2)
which gives a compatibility condition.
Definition 1.
We say that $\varphi:\overline{\Lambda}\times\mathbb{R}\to\mathbb{R}$ is a moving flock if
$\varphi(l,t)=\bar{\varphi}(l)+\bar{v}t\quad\text{ for }(l,t)\in\overline{\Lambda}\times\mathbb{R},$
where $\bar{v}=\pm\sqrt{\frac{\alpha}{\beta}}$. We say that $\varphi:\overline{\Lambda}\times\mathbb{R}\to\mathbb{R}$ is a stationary flock if
$\varphi(l,t)=\bar{\varphi}(l)\quad\text{ for }(l,t)\in\overline{\Lambda}\times\mathbb{R}.$
By symmetry, we only need to consider the moving flocks with $\bar{v}=\sqrt{\alpha/\beta}$. The main questions of interest in our paper are as follows.
  • (1)
    Let $\bar{v}=\sqrt{\alpha/\beta}$ and $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Can we find a control $u\in\mathcal{A}_{M}$ and an associated $\varphi_{1}$ such that $w=\varphi_{1}$ for all $t\geq T$, where $T>0$ is sufficiently large and $\varphi_{1}$ is a moving flock?
  • (2)
    Let $\bar{v}=\sqrt{\alpha/\beta}$ and $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Take an arbitrary moving flock $\varphi_{2}(l,t)=\bar{\varphi}_{2}(l)+\bar{v}t$. Can we find a control $u\in\mathcal{A}_{M}$ such that $w-\varphi_{2}$ is constant for all $t\geq T$ where $T>0$ is sufficiently large?
  • (3)
    Let $\bar{v}=0$ and $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Take an arbitrary stationary flock $\varphi_{3}(l,t)=\bar{\varphi}_{3}(l)$. Can we find a control $u\in\mathcal{A}_{M}$ such that $w-\varphi_{3}$ is constant for all $t\geq T$ where $T>0$ is sufficiently large?
1.3. Main results
Write $x_{l}(t)=w(l,t)$ and $v_{l}(t)=w_{t}(l,t)$ for $l\in\Lambda$ and $t\geq 0$. Then, (1.1) becomes a system of ODEs, for $(l,t)\in\Lambda\times(0,\infty)$,
$\begin{cases}\dot{x}_{l}(t)=v_{l}(t),\\ \displaystyle\dot{v}_{l}(t)=\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}+(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)+u_{l}(t),\end{cases}$ (1.3)
with the given initial condition
$\begin{cases}x_{l}(0)=w_{0}(l),\\ v_{l}(0)=w_{1}(l).\end{cases}$
Assumption 1.
We will assume that $M$ in (1.2) satisfies
$M>\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}.$ (1.4)
Theorem 1.1.
Assume (1.4). Let $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Then, there exist $u\in\mathcal{A}_{M}$ and a moving flock
$\varphi_{1}(l,t)=\bar{x}_{1l}+(t-T)\bar{v}=\bar{x}_{1l}-\bar{v}T+\bar{v}t\quad\text{ for }(l,t)\in\overline{\Lambda}\times\mathbb{R}.$
where $\bar{x}_{1l}\in\mathbb{R}$ are constants such that $w=\varphi_{1}$ for all $t\geq T$ where $T$ is taken to be sufficiently large.
Theorem 1.2.
Assume (1.4). Let $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Let $\varphi_{2}(l,t)=\bar{\varphi}_{2}(l)+\bar{v}t$ be an arbitrary flock, where $\bar{\varphi}_{2}$ satisfies
$|-\Delta\bar{\varphi}_{2}(l)|<M\quad\text{ for }l\in\overline{\Lambda}$ (1.5)
Then, there exist $u\in\mathcal{A}_{M}$ and a constant $a\in\mathbb{R}$ such that the solution to (1.3) satisfies
$\begin{cases}x_{l}(t)=\bar{\varphi}_{2}(l)+a+\bar{v}t,\\ v_{l}(t)=\bar{v}=\sqrt{\alpha/\beta}\end{cases}$
for $l\in\overline{\Lambda}$ and $t>T$. In particular, we have $w=\varphi_{2}+a$ for all $t\geq T$ where $T$ is taken to be sufficiently large.
As a by-product of the proofs of Theorems 1.11.2, we have the following result.
Theorem 1.3.
Assume (1.4). Let $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Let $\varphi_{3}(l,t)=\bar{\varphi}_{3}(l)$ be an arbitrary stationary flock, where $\bar{\varphi}_{3}$ satisfies
$|-\Delta\bar{\varphi}_{3}(l)|<M\quad\text{ for }l\in\overline{\Lambda}.$ (1.6)
Then, there exist $u\in\mathcal{A}_{M}$ and a constant $a\in\mathbb{R}$ such that the solution to (1.3) satisfies
$\begin{cases}x_{l}(t)=\bar{\varphi}_{3}(l)+a,\\ v_{l}(t)=\bar{v}=0.\end{cases}$
for $l\in\overline{\Lambda}$ and $t>T$. In particular, we have $w=\varphi_{3}+a$ for all $t\geq T$ where $T$ is taken to be sufficiently large.
We end the introduction by noting that in recent years, there have been important advances in control theory for kinetic models, particularly the control to flocking of the Cucker–Smale model [14], which serves as our source of inspiration for studying the problem of controlling the propagation of lattice/chains nonlinear waves. We note that the assumption (1.4) on $M$ in Theorems 1.11.3 is essentially optimal: see Remark 2.
Organization of the paper
The paper is organized as follows. In Section 2, we prove Theorem 1.1. The proofs of Theorems 1.21.3 are given in Section 3. Finally, section 4 contains a connection between the flocking problem and a minimal time problem in the framework of optimal control theory and Hamilton–Jacobi equations.
2. Proof of Theorem 1.1
Our strategy to prove Theorem 1.1 is as follows. We start by using a control $u$ that drives the system toward zero velocity. To do so, we first design $u$ so that the particles’ velocities become sufficiently small after some finite time. Once the system is close to zero velocity, we can find a control in $\mathcal{A}_{M}$ that brings it to exactly zero velocity, a stationary flock, after another finite amount of time. Finally, we apply a third type of control to steer the system from a stationary flock to a moving flock with the desired group velocity $\bar{v}=\sqrt{\alpha/\beta}$. It is important to note that all the controls designed are explicit in terms of $(x_{l}(t))_{l\in\overline{\Lambda}}$ and $(v_{l}(t))_{l\in\overline{\Lambda}}$.
In the process of driving the system to zero velocity, we will study the Lyapunov function
$V(t)=\frac{1}{2}\sum_{l\in\overline{\Lambda}}\left[|v_{l}(t)|^{2}+\frac{1}{2}\sum_{l^{\prime}\sim l}\frac{(x_{l^{\prime}}(t)-x_{l}(t))^{2}}{h^{2}}\right].$
It is immediate that $V\geq 0$. We will now prove
$\dot{V}(t)=\sum_{l\in\overline{\Lambda}}\left[(\alpha-\beta|v_{l}(t)|^{2})|v_{l}(t)|^{2}+(u_{l}(t),v_{l}(t))\right]$
which shows that by choosing $u_{l}(t)$, we can control $\dot{V}$. For clarity in our writing, we use the usual way to write dot product $(a,b)=ab$ for $a,b\in\mathbb{R}$.
We have
$\dot{V}(t)$ $=\sum_{l\in\overline{\Lambda}}\left[(\dot{v}_{l}(t),v_{l}(t))+\frac{1}{2}\sum_{l^{\prime}\sim l}\left(\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}},v_{l^{\prime}}(t)-v_{l}(t)\right)\right]$
$=\sum_{l\in\overline{\Lambda}}\Bigg[\left(\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}+(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)+u_{l}(t),v_{l}(t)\right)$
$\qquad\qquad\qquad+\frac{1}{2}\left(\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}},-v_{l}(t)\right)+\frac{1}{2}\sum_{l^{\prime}\sim l}\left(\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}},v_{l^{\prime}}(t)\right)\Bigg]$
$=\sum_{l\in\overline{\Lambda}}\left((\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)+u_{l}(t),v_{l}(t)\right)$
$=\sum_{l\in\overline{\Lambda}}\left[(\alpha-\beta|v_{l}(t)|^{2})|v_{l}(t)|^{2}+(u_{l}(t),v_{l}(t))\right].$
Let $\gamma>\max\left\{1,\frac{1}{M}\sqrt{\frac{\alpha^{3}}{\beta}}\right\}$ be fixed and define
$a_{1}=\frac{1}{\gamma}\sqrt{\frac{\alpha}{\beta}}\qquad\text{ and }\qquad a_{2}=\gamma\sqrt{\frac{\alpha}{\beta}}.$
When $t\in[0,T_{0}]$ for some $T_{0}$ to be defined later, we will use the control $u_{l}$ based on $v_{l}$ in the following way:
$u_{l}(v_{l})=\begin{cases}0&\text{if }|v_{l}|\geq 2a_{2},\\ -M\frac{v_{l}}{|v_{l}|}\left(2-\frac{|v_{l}|}{a_{2}}\right)&\text{if }|v_{l}|\in(a_{2},2a_{2}),\\ -M\frac{v_{l}}{|v_{l}|}&\text{if }|v_{l}|\in[a_{1},a_{2}],\\ -M\frac{v_{l}}{a_{1}}&\text{if }|v_{l}|<a_{1}.\end{cases}$ (2.1)
Of course, $u_{l}$ is Lipschitz in $v_{l}$. See Remark 1 for a discussion on this choice of control. For each time $t$, we can split the indices $l$ into 3 sets:
$\begin{cases}I_{2}(t)=\{l\in\overline{\Lambda}\,:\,|v_{l}(t)|>a_{2}\},\\ I_{1}(t)=\{l\in\overline{\Lambda}\,:\,|v_{l}(t)|\in[a_{1},a_{2}]\},\\ I_{0}(t)=\{l\in\overline{\Lambda}\,:\,|v_{l}(t)|<a_{1}\}.\end{cases}$
For $l\in I_{2}(t)$, $u_{l}v_{l}=0$ if $|v_{l}|\geq 2a_{2}$ or
$u_{l}v_{l}=-M\frac{v_{l}}{|v_{l}|}\left(2-\frac{|v_{l}|}{a_{2}}\right)v_{l}=-M|v_{l}|\left(2-\frac{|v_{l}|}{a_{2}}\right)<0$
if $|v_{l}|\in(a_{2},2a_{2})$. As $\gamma>1$,
$(\alpha-\beta|v_{l}(t)|^{2})|v_{l}(t)|^{2}<(\alpha-\beta a_{2}^{2})|v_{l}(t)|^{2}<(\alpha-\gamma^{2}\alpha)|v_{l}(t)|^{2}<0.$
Hence,
$(\alpha-\beta|v_{l}(t)|^{2})|v_{l}(t)|^{2}+u_{l}v_{l}<(\alpha-\gamma^{2}\alpha)a_{2}^{2}.$
We define
$C_{2}:=(\gamma^{2}-1)\alpha a_{2}^{2}>0.$
For $l\in I_{1}(t)$, we see $u_{l}v_{l}=-M|v_{l}|<0$. Also,
$(\alpha-\beta|v_{l}|^{2})|v_{l}|^{2}+u_{l}v_{l}=|v_{l}|(\alpha|v_{l}|-\beta|v_{l}|^{3}-M).$ (2.2)
Consider the function $f(x)=\alpha x-\beta x^{3}$. It is clear that
$\max_{x\geq 0}f(x)=f\left(\sqrt{\frac{\alpha}{3\beta}}\right)=\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}<M$ (2.3)
by our assumption (1.4). Applying this to (2.2), we see
$(\alpha-\beta|v_{l}|^{2})|v_{l}|^{2}+u_{l}v_{l}<|v_{l}|\left(\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}-M\right)<a_{1}\left(\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}-M\right).$ (2.4)
We define
$C_{1}:=a_{1}\left(M-\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}\right)>0.$
Finally, for $l\in I_{0}(t)$, $u_{l}v_{l}=-M\frac{|v_{l}|^{2}}{a_{1}}$. Since $\gamma>\frac{1}{M}\sqrt{\frac{\alpha^{3}}{\beta}}$, we have $a_{1}<\frac{M}{\alpha}$. Thus,
$(\alpha-\beta|v_{l}|^{2})|v_{l}|^{2}+u_{l}v_{l}=|v_{l}|^{2}\left(\alpha-\beta|v_{l}|^{2}-\frac{M}{a_{1}}\right)\leq-\beta|v_{l}|^{4}.$
Therefore, we have shown that when $u_{l}$ is as described in (2.1),
$\dot{V}(t)\leq\sum_{l\in I_{2}(t)}-C_{2}+\sum_{l\in I_{1}(t)}-C_{1}+\sum_{l\in I_{0}(t)}-\beta|v_{l}(t)|^{4}.$ (2.5)
Thus, $V(t)$ is nonincreasing and, in particular, $V(t)$ is bounded from above by $V(0)$. Since both terms in $V$ are nonnegative, this means they must both be bounded from above. Additionally, for $l\in\overline{\Lambda}$,
$\sum_{l^{\prime}\sim l}\frac{(x_{l^{\prime}}(t)-x_{l}(t))^{2}}{h^{2}}\leq V(0),$
which implies that
$\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}\right|\leq C.$ (2.6)
Therefore, we have
$|\dot{v}_{l}(t)|\leq C(1+\alpha+|u_{l}(t)|)\leq C\quad\text{or}\quad|\dot{v}_{l}(t)|\leq C(1+\beta C^{2}+|u_{l}(t)|)\leq C.$
Here, $C=C(V(0),h,\alpha,\beta)>0$.
Lemma 2.1.
The system (1.3) satisfies
$\lim_{t\to\infty}v_{l}(t)=0\quad\text{ for all $l\in\overline{\Lambda}$}$
when $u_{l}$ is given by (2.1).
Proof.
Assume this is not the case. Then there exists a $c_{0}>0$, an index $l$, and a sequence of times $t_{k}\to\infty$ such that
$|v_{l}(t_{k})|>c_{0}.$
Since $|\dot{v}_{l}(t)|\leq C$, there is a $\tau>0$ such that
$|v_{l}(t)|>\frac{c_{0}}{2}\quad\text{ for all $t\in(t_{k}-\tau,t_{k}+\tau)$, $k\in\mathbb{N}$.}$
Then
$\dot{V}(t)<\max\left\{-C_{2},-C_{1},-\beta\frac{c_{0}^{4}}{16}\right\}\quad\text{ for all $t\in(t_{k}-\tau,t_{k}+\tau)$, $k\in\mathbb{N}$.}$
This shows that
$\lim_{t\to\infty}V(t)=-\infty,$
which contradicts the fact that $V\geq 0$. Thus, $\lim_{t\to\infty}v_{l}(t)=0$. ∎
Remark 1.
In (2.1) above, we define $u_{l}=u_{l}(v_{l})$ in such a way that we do not need to use the control $u_{l}$ when $|v_{l}|\geq 2a_{2}$. This choice is not necessary, but for practical purposes, having the control set to zero might be preferred, and allowing the system to be driven by the large velocities in the system might be preferable. Another way, which is simpler but requires a higher magnitude of control, is to define $u_{l}$ as
$u_{l}(v_{l})=\begin{cases}-M\frac{v_{l}}{|v_{l}|}&\text{if }|v_{l}|\geq a_{1},\\ -M\frac{v_{l}}{a_{1}}&\text{if }|v_{l}|<a_{1}.\end{cases}$
The proof that this control drives the system towards zero velocity would be identical to the proof above except for the definition of $I_{1}(t)$.
Remark 2.
We note that (2.2) and (2.4) show the need for the assumption on $M$. The magnitude of $u$ has to be bounded by $M$. However, if we would like $\dot{V}<0$, then $u$ must have a large enough magnitude to cancel out $(\alpha-\beta|v_{l}|^{2})|v_{l}|$ which is equal to $\frac{2}{3\sqrt{2}}\sqrt{\frac{\alpha^{3}}{\beta}}$ when $|v_{l}|=\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}$. Thus, $M>\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}$ ensures $\dot{V}<0$. If $t\mapsto V(t)$ is not decreasing, Lemma 2.1 might not hold true.
Lemma $2.$
The system (1.3) satisfies
$\lim_{t\to\infty}\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}=0\quad\text{ for all $l\in\overline{\Lambda}$}$ (2.7)
when $u_{l}$ is given by (2.1).
Proof.
Since $|\dot{x}_{l}(t)|=|v_{l}(t)|\leq\sqrt{2V(0)}$, the functions $x_{l}(t)$ are Lipschitz with constant $L:=\sqrt{2V(0)}$. Then $x_{l^{\prime}}(t)-x_{l}(t)$ are Lipschitz with Lipschitz constants less than or equal to $2L$. Assume for the sake of contradiction that (2.7) does not hold. In light of (2.6), there exists an $l$ and a sequence $\{t_{k}\}$ such that $t_{k}\to\infty$ and
$\lim_{k\to\infty}\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t_{k})-x_{l}(t_{k})}{h^{2}}=c\neq 0.$ (2.8)
There exists a constant $K>0$ such that when $k>K$
$\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}-c\right|$
$=\,$ $\left|\left[\sum_{l^{\prime}\sim l}\frac{(x_{l^{\prime}}(t)-x_{l}(t))-((x_{l^{\prime}}(t_{k})-x_{l}(t_{k}))+(x_{l^{\prime}}(t_{k})-x_{l}(t_{k}))}{h^{2}}\right]-c\right|$
$\leq\,$ $\frac{8nL}{h^{2}}|t_{k}-t|+\frac{|c|}{8}.$
Additionally, there exists a sequence $(\eta_{k})_{k\geq K}$ such that when $t>t_{k}$,
$\begin{cases}|v_{l}(t)|<\eta_{k}<a_{1},\\ |(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)|<\eta_{k}<a_{1}\end{cases}$ (2.9)
and $\eta_{k}\to 0$ due to Lemma 2.1. Thus,
$\begin{split}&|\dot{v}_{l}(t)-c|=\,\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}-c+(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)+u_{l}(t)\right|\\ \leq\,&\frac{8nL}{h^{2}}|t_{k}-t|+\frac{|c|}{8}+\eta_{k}+\frac{M}{a_{1}}\eta_{k}.\end{split}$ (2.10)
Let $\tau=\min\{1,\frac{|c|h^{2}}{16nL}\}$. Using (LABEL:31), we obtain
$|v_{l}(t_{k}+\tau)-c\tau|-|v_{l}(t_{k})|\leq|v_{l}(t_{k}+\tau)-v_{l}(t_{k})-c\tau|$
$=\,$ $\left|\int^{t_{k}+\tau}_{t_{k}}\left(\dot{v}_{l}(s)-c\right)\,ds\right|\leq\,\int^{t_{k}+\tau}_{t_{k}}|\dot{v}_{l}(s)-c|\,ds$
$\leq$ $\tau\left(\frac{8nL\tau}{h^{2}}+\frac{|c|}{8}+\eta_{k}+\frac{M}{a_{1}}\eta_{k}\right)\leq\tau\left(\frac{|c|}{2}+\frac{|c|}{8}+\eta_{k}+\frac{M}{a_{1}}\eta_{k}\right).$
Thus,
$|v_{l}(t_{k}+\tau)-c\tau|\leq\,|v_{l}(t_{k})|+\tau\left(\frac{|c|}{2}+\frac{|c|}{8}+\eta_{k}+\frac{M}{a_{1}}\eta_{k}\right)$
$\leq\,$ $\eta_{k}+\tau\left(\frac{|c|}{2}+\frac{|c|}{8}+\eta_{k}+\frac{M}{a_{1}}\eta_{k}\right).$
Using the reverse triangle inequality again, we have
$|c\tau|-|v_{l}(t_{k}+\tau)|\leq\eta_{k}+\frac{|c|\tau}{2}+\frac{|c|\tau}{8}+\left(1+\frac{M}{a_{1}}\right)\tau\eta_{k},$
so
$\frac{3|c|\tau}{8}-\tau\eta_{k}\left(1+\frac{M}{a_{1}}\right)-\eta_{k}\leq|v_{l}(t_{k}+\tau)|\quad\text{ for all }k>K.$ (2.11)
Since $\eta_{k}\to 0$, we know there exists $K_{2}$ such that when $k>K_{2}$, $\eta_{k}(1+\frac{M}{a_{1}})\leq\frac{|c|}{16}$ and $\eta_{k}<\frac{|c|\tau}{16}$. Thus, when $k>\max\{K,K_{2}\}$,
$|v_{l}(t_{k}+\tau)|\geq\frac{|c|\tau}{4}.$
However, if we take the limit as $k\to\infty$, we have $\frac{|c|\tau}{4}\leq 0$. Therefore, $c=0$, which contradicts (2.8). The proof is complete. ∎
Therefore, there exists a $T_{0}$ such that if (2.1) is used for $0<t\leq T_{0}$,
$|v_{l}(T_{0})|<\epsilon\quad\text{ and }\quad\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{0})-x_{l}(T_{0})}{h^{2}}\right|<\epsilon$
for some $\epsilon$ satisfying
$0<\epsilon<\min\left\{1,\frac{M}{2n+1+\frac{8n}{h^{2}}+2\alpha+8\beta}\right\}.$ (2.12)
It is important to note that $\epsilon$ can be chosen explicitly here with the above bound, which is useful in practice. We are now ready to finish the proof of Theorem 1.1.
Proof of Theorem 1.1.
Define $T_{1,l}=T_{0}+\frac{1}{\epsilon}|v_{l}(T_{0})|$ for each $l\in\overline{\Lambda}$ and $T_{1}=\max_{l\in\overline{\Lambda}}T_{1,l}$. For $T_{0}<t\leq T_{1}$, we define
$u_{l}(t)=\dot{v}_{l}(t)-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}-(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)$ (2.13)
where $v_{l}(t)$ satisfies the following:
$v_{l}(t)=\begin{cases}v_{l}(T_{0})-\epsilon(t-T_{0})\frac{v_{l}(T_{0})}{|v_{l}(T_{0})|}\quad&\text{for }t\in[T_{0},T_{1,l}],\\ 0&\text{for }t>T_{1,l}.\end{cases}$ (2.14)
and
$x_{l}(t)$ $=x_{l}(T_{0})+\int_{T_{0}}^{t}\dot{x}_{l}(s)\,ds=x_{l}(T_{0})+\int_{T_{0}}^{t}\left(v_{l}(T_{0})-\epsilon(s-T_{0})\frac{v_{l}(T_{0})}{|v_{l}(T_{0})|}\right)\,ds$
$=x_{l}(T_{0})+v_{l}(T_{0})(t-T_{0})-\frac{\epsilon}{2}\frac{v_{l}(T_{0})}{|v_{l}(T_{0})|}(t-T_{0})^{2}$
when $t\in[T_{0},T_{1,l}]$ and $x_{l}(t)=x_{l}(T_{1,l})$ when $t>T_{1,l}$. Notice
$|x_{l}(t)-x_{l}(T_{0})|$ $\leq|v_{l}(T_{0})|(T_{1,l}-T_{0})+\frac{\epsilon}{2}(T_{1,l}-T_{0})^{2}$
$\leq\frac{|v_{l}(T_{0})|^{2}}{\epsilon}+\frac{\epsilon}{2}\frac{|v_{l}(T_{0})|^{2}}{\epsilon^{2}}<2\epsilon.$
We must show that $|u_{l}|<M$. We first observe that when $t\in(T_{1,l},T_{1}]$,
$u_{l}(t)=0-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1,l})-x_{l}(T_{1,l})}{h^{2}}-0=-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1,l})-x_{l}(T_{1,l})}{h^{2}}.$
Thus,
$|u_{l}(t)|$ $=\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1,l})-x_{l}(T_{1,l})}{h^{2}}\right|$
$=\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1,l})-x_{l^{\prime}}(T_{0})+x_{l^{\prime}}(T_{0})-x_{l}(T_{0})+x_{l}(T_{0})-x_{l}(T_{1,l})}{h^{2}}\right|$
$\leq\sum_{l^{\prime}\sim l}\left(\frac{2\epsilon}{h^{2}}+\epsilon+\frac{2\epsilon}{h^{2}}\right)=\frac{8n\epsilon}{h^{2}}+2n\epsilon$
when $t\in(T_{1,l},T_{1}].$ Similarly, when $t\in(T_{0},T_{1,l}]$,
$u_{l}(t)$ $=-\epsilon\frac{v_{l}(T_{0})}{|v_{l}(T_{0})|}-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}$
$\qquad-\left(\alpha-\beta\left|v_{l}(T_{0})-\epsilon(t-T_{0})\frac{v_{l}(T_{0})}{|v_{l}(T_{0})|}\right|^{2}\right)\left(v_{l}(T_{0})-\epsilon(t-T_{0})\frac{v_{l}(T_{0})}{|v_{l}(T_{0})|}\right).$
Hence, using the bound on $\epsilon$ given in (2.12) we see
$|u_{l}(t)|$
$\leq\,$ $\epsilon+\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}\right|+\alpha\left|v_{l}(T_{0})\right|+\alpha\left|\epsilon(T_{1,l}-T_{0})\right|+\beta\left(\left|v_{l}(T_{0})\right|+\epsilon|T_{1,l}-T_{0}|\right)^{3}$
$\leq\,$ $\epsilon+\frac{8n\epsilon}{h^{2}}+2n\epsilon+\alpha\epsilon+\alpha\epsilon+\beta\left(\left|v_{l}(T_{0})\right|+\epsilon\right)^{3}$
$=\,$ $\epsilon\left(2n+1+\frac{8n}{h^{2}}+2\alpha+8\beta\epsilon^{2}\right)<M.$
Thus, we have designed a control that drives the system to zero velocity at $T_{1}$. The system is now at a stationary flock at $t=T_{1}$.
Finally, let
$u_{l}(t)=-(\alpha-\beta|v_{l}|^{2})v_{l}-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}+w\quad\text{ for $t\in[T_{1},T_{2}]$,}$
where
$T_{2}=T_{1}+\frac{2\sqrt{\alpha}}{M\sqrt{\beta}}\quad\text{ and }\quad w=\frac{\bar{v}}{T_{2}-T_{1}}=\frac{M}{2}.$
We notice that the first terms of $u_{l}$ will cancel out the other terms on the right side of the second equation in (1.3) so that
$\dot{v}_{l}(t)=w=\frac{M}{2}.$
Additionally, $v_{l}(T_{1})=0$. Thus, $v_{l}(t)$ is the same for all $l\in\overline{\Lambda}$ and
$v_{l}(t)=\frac{M(t-T_{1})}{2}\quad\text{ for }t\in[T_{1},T_{2}].$
Therefore,
$\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}=\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1})-x_{l}(T_{1})}{h^{2}}\quad\text{ for all $t\in[T_{1},T_{2}]$.}$
Recall that this term has magnitude less than $\epsilon\ll M/2$. With this information, we can show that $|u_{l}|<M$. As $0\leq v_{l}(t)\leq\bar{v}$ we can see using (2.3) that
$0\leq(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)\leq\frac{2}{3\sqrt{3}}\sqrt{\frac{\alpha^{3}}{\beta}}<M.$
Thus,
$|u_{l}(t)|\leq\left|-(\alpha-\beta|v_{l}|^{2})v_{l}+\frac{M}{2}\right|+\left|-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}\right|\leq\frac{M}{2}+\epsilon<M.$
At $t=T_{2}$, we get
$v_{l}(T_{2})=\bar{v}=\sqrt{\frac{\alpha}{\beta}},$
and
$\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{2})-x_{l}(T_{2})}{h^{2}}\right|=\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1})-x_{l}(T_{1})}{h^{2}}\right|\leq\epsilon.$
For $t\geq T_{2}$, we simply keep $v_{l}(t)=v_{l}(T_{2})=\bar{v}$ for all $l\in\overline{\Lambda}$. Then, we get
$x_{l}(t)=x_{l}(T_{2})+\bar{v}(t-T_{2})\quad\text{ for all }t\geq T_{2}.$
The proof is complete with $T=T_{2}$ and $\bar{x}_{1l}=x_{l}(T_{2})$ for $l\in\overline{\Lambda}$. ∎
Remark 3.
It is important to note that, in the above proof, for $l\in\overline{\Lambda}$,
$\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{2})-x_{l}(T_{2})}{h^{2}}\right|=\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1})-x_{l}(T_{1})}{h^{2}}\right|\leq\epsilon,$
where $\epsilon>0$ can be chosen as small as we want. This is an important point that we will need to use in the proof of Theorem 1.2.
3. Proofs of Theorems 1.21.3
3.1. Proof of Theorem 1.2
Our strategy to prove Theorem 1.2 is as follows. By Theorem 1.1, we have driven the system to a moving flock at time $T_{2}$. By designing an explicit control, we drive the system to a moving flock where every particle is at the same position at $T_{3}>T_{2}$. From this intermediate moving flock where all particles are at the same position, we will finally design another explicit control driving the system to the desired flock at $T_{4}>T_{3}$. Along the way, we need to make sure that the changes of the positions of the particles and their corresponding velocities are sufficiently small so that the controls always satisfy the given bound $M$.
Proof of Theorem 1.2.
By the proof of Theorem 1.1, we know that given any initial conditions, we can drive the system to a moving flock $\varphi_{1}(l,t)=\bar{\varphi}_{1}(l)+\bar{v}t$ at time $t=T_{2}$ satisfying
$\left|\sum_{l^{\prime}\sim l}\frac{\bar{\varphi}_{1}(l^{\prime})-\bar{\varphi}_{1}(l)}{h^{2}}\right|\leq\epsilon$ (3.1)
for all $l\in\overline{\Lambda}$. We will then adjust the control to drive the system to
$\begin{cases}x_{l}(T_{3})=x_{j}(T_{3}),\\ v_{l}(T_{3})=v_{j}(T_{3})=\bar{v},\end{cases}$
for all $l,j\in\overline{\Lambda}$. That is, the system is at a moving flock where every particle is at the same position at $T_{3}>T_{2}$. Let
$\bar{y}=\frac{1}{D^{n}}\sum_{l\in\overline{\Lambda}}\bar{\varphi}_{1}(l),$
and let $T_{3}>T_{2}$ satisfy
$\begin{split}&\frac{3\alpha|\bar{y}-\bar{\varphi}_{1}(l)|(T_{3}-T_{2})^{2}+6.75\sqrt{\alpha\beta}|\bar{y}-\bar{\varphi}_{1}(l)|^{2}(T_{3}-T_{2})}{(T_{3}-T_{2})^{3}}\\ &+\frac{3.375\beta|\bar{y}-\bar{\varphi}_{1}(l)|^{3}+6|\bar{y}-\bar{\varphi}_{1}(l)|(T_{3}+T_{2})}{(T_{3}-T_{2})^{3}}<M-\epsilon\end{split}$ (3.2)
for all $l\in\overline{\Lambda}$.
Remark 4.
We note that our choice of $\bar{y}$ is reasonable, but another point could be chosen. In fact, if we choose $\bar{y}$ such that it minimizes $\max_{l\in\Lambda}|y-\bar{\varphi}_{1}(l)|$ then we can reduce $T_{3}$.
We then set
$\begin{split}u_{l}(t)=&-\left[1+\frac{2(t-T_{2})^{3}}{(T_{3}-T_{2})^{3}}-\frac{3(t-T_{2})^{2}}{(T_{3}-T_{2})^{2}}\right]\sum_{l^{\prime}\sim l}\frac{\bar{\varphi_{1}}(l^{\prime})-\bar{\varphi_{1}}(l)}{h^{2}}\\ &+\frac{12\alpha}{(T_{3}-T_{2})^{3}}(\bar{y}-\bar{\varphi}_{1}(l))(t-T_{2})(T_{3}-t)\\ &+\frac{108\sqrt{\alpha\beta}}{(T_{3}-T_{2})^{6}}(\bar{y}-\bar{\varphi}_{1}(l))^{2}(t-T_{2})^{2}(T_{3}-t)^{2}\\ &+\frac{216\beta}{(T_{3}-T_{2})^{9}}(\bar{y}-\bar{\varphi}_{1}(l))^{3}(t-T_{2})^{3}(T_{3}-t)^{3}+\frac{6(\bar{y}-\bar{\varphi_{1}}(l))}{(T_{3}-T_{2})^{3}}(T_{3}+T_{2}-2t)\end{split}$
for $t\in[T_{2},T_{3}]$. With this choice of control and the fact that $v(T_{2})=\bar{v}$, we find
$v_{l}(t)=\bar{v}+\frac{6(\bar{y}-\bar{\varphi_{1}}(l))}{(T_{3}-T_{2})^{3}}(t-T_{2})(T_{3}-t)$
and
$x_{l}(t)=\bar{\varphi}_{1}(l)+(t-T_{2})\bar{v}+\frac{3(t-T_{2})^{2}}{(T_{3}-T_{2})^{2}}(\bar{y}-\bar{\varphi_{1}}(l))-\frac{2(t-T_{2})^{3}}{(T_{3}-T_{2})^{3}}(\bar{y}-\bar{\varphi_{1}}(l)).$
Since $x_{l}(T_{3})=\bar{y}+(T_{3}-T_{2})\bar{v}$, we deduce $x_{l}(T_{3})=x_{j}(T_{3})$ and $v_{l}(T_{3})=v_{j}(T_{3})=\bar{v}$ for all $l,j\in\overline{\Lambda}$. However, we must show that $|u_{l}(t)|<M$. Notice that $(t-T_{2})(T_{3}-t)$ has a maximum value of $\frac{(T_{3}-T_{2})^{2}}{4}$ and
$1+\frac{2(t-T_{2})^{3}}{(T_{3}-T_{2})^{3}}-\frac{3(t-T_{2})^{2}}{(T_{3}-T_{2})^{2}}\leq 1\quad\text{ for $t\in[T_{2},T_{3}]$.}$
Combining these facts with (3.1) and (LABEL:T2cond) shows
$|u_{l}(t)|$
$\leq\,$ $\left|\sum_{l^{\prime}\sim l}\frac{\bar{\varphi}_{1}(l^{\prime})-\bar{\varphi}_{1}(l)}{h^{2}}\right|+\frac{12\alpha}{(T_{3}-T_{2})^{3}}|\bar{y}-\bar{\varphi}_{1}(l)|\frac{(T_{3}-T_{2})^{2}}{4}$
$\,+\frac{108\sqrt{\alpha\beta}}{(T_{3}-T_{2})^{6}}|\bar{y}-\bar{\varphi}_{1}(l)|^{2}\frac{(T_{3}-T_{2})^{4}}{16}+\frac{216\beta}{(T_{3}-T_{2})^{9}}|\bar{y}-\bar{\varphi}_{1}(l)|^{3}\frac{(T_{3}-T_{2})^{6}}{64}$
$\,+\frac{6(\bar{y}-\bar{\varphi}_{1}(l))}{(T_{3}-T_{2})^{3}}(T_{3}+T_{2})$
$\leq\,$ $\epsilon+\frac{3\alpha|\bar{y}-\bar{\varphi}_{1}(l)|(T_{3}-T_{2})^{2}+6.75\sqrt{\alpha\beta}|\bar{y}-\bar{\varphi}_{1}(l)|^{2}(T_{3}-T_{2})}{(T_{3}-T_{2})^{3}}$
$\,\,+\frac{3.375\beta|\bar{y}-\bar{\varphi}_{1}(l)|^{3}+6|\bar{y}-\bar{\varphi}_{1}(l)|(T_{3}+T_{2})}{(T_{3}-T_{2})^{3}}<M.$
From the intermediate moving flock where all particles are at the same position, we will finally design the control for $t\in[T_{3},T_{4}]$ that drives the system to the desired flock
$x_{l}(t)=\bar{\varphi}_{2}(l)+a+\bar{v}t\quad\text{ for }l\in\overline{\Lambda},t\geq T_{4}.$
By the assumptions, $|-\Delta\bar{\varphi}_{2}|<M$. Let $T_{4}>T_{3}$ satisfy
$\begin{split}&\frac{3\alpha|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|(T_{4}-T_{3})^{2}+6.75\sqrt{\alpha\beta}|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|^{2}(T_{4}-T_{3})}{(T_{4}-T_{3})^{3}}\\ &+\frac{3.375\beta|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|^{3}+6|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|(T_{4}+T_{3})}{(T_{4}-T_{3})^{3}}<M-|\Delta\bar{\varphi}_{2}(l)|\end{split}$ (3.3)
for all $l\in\overline{\Lambda}$.
Now from $T_{3}$ to $T_{4}$, letting
$\begin{split}u_{l}(t)=&\left[-\frac{3(t-T_{3})^{2}}{(T_{4}-T_{3})^{2}}+\frac{2(t-T_{3})^{3}}{(T_{4}-T_{3})^{3}}\right]\sum_{l^{\prime}\sim l}\frac{\bar{\varphi}_{2}(l^{\prime})-\bar{\varphi}_{2}(l)}{h^{2}}\\ &+\frac{12\alpha}{(T_{4}-T_{3})^{3}}(\bar{\varphi}_{2}(l)-x_{l}(T_{3}))(t-T_{3})(T_{4}-t)\\ &+\frac{108\sqrt{\alpha\beta}}{(T_{4}-T_{3})^{6}}(\bar{\varphi}_{2}(l)-x_{l}(T_{3}))^{2}(t-T_{3})^{2}(T_{4}-t)^{2}\\ &+\frac{216\beta}{(T_{4}-T_{3})^{9}}(\bar{\varphi}_{2}(l)-x_{l}(T_{3}))^{3}(t-T_{3})^{3}(T_{4}-t)^{3}\\ &+\frac{6(\bar{\varphi}_{2}(l)-x_{l}(T_{3}))}{(T_{4}-T_{3})^{3}}(T_{4}+T_{3}-2t),\end{split}$
we arrive at
$v_{l}(t)=\bar{v}+\frac{6(\bar{\varphi}_{2}(l)-x_{l}(T_{3}))}{(T_{4}-T_{3})^{3}}(t-T_{3})(T_{4}-t)$
and
$x_{l}(t)=x_{l}(T_{3})+(t-T_{3})\bar{v}+\frac{3(t-T_{3})^{2}}{(T_{4}-T_{3})^{2}}(\bar{\varphi}_{2}(l)-x_{l}(T_{3}))-\frac{2(t-T_{3})^{3}}{(T_{4}-T_{3})^{3}}(\bar{\varphi}_{2}(l)-x_{l}(T_{3})).$
Then,
$\begin{cases}x_{l}(T_{4})=(T_{4}-T_{3})\bar{v}+\bar{\varphi}_{2}(l),\\ v_{l}(T_{4})=\bar{v}.\end{cases}$
Once again, we would like to ensure $|u_{l}(t)|<M$. For $l\in\overline{\Lambda}$, we estimate
$|u_{l}(t)|$
$\leq\,$ $|\Delta\bar{\varphi}_{2}(l)|+\frac{3\alpha|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|(T_{4}-T_{3})^{2}+6.75\sqrt{\alpha\beta}|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|^{2}(T_{4}-T_{3})}{(T_{4}-T_{3})^{3}}$
$+\frac{3.375\beta|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|^{3}+6|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|(T_{4}+T_{3})}{(T_{4}-T_{3})^{3}}<M.$
Finally, for $t>T_{4}$, we will let
$u_{l}(t)=-\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}},$
so $v_{l}(t)=\bar{v}$, and $x_{l}(t)=(t-T_{3})\bar{v}+\bar{\varphi}_{2}(l)$ for $t\geq T_{4}$. Here, $a=-T_{3}\bar{v}$. ∎
Remark 5.
We note that assumption (1.5) in Theorem 1.2 can be relaxed to either
$M\leq-\Delta\bar{\varphi}_{2}(l)<M\quad\text{ for all }l\in\overline{\Lambda}$ (3.4)
or
$M<-\Delta\bar{\varphi}_{2}(l)\leq M\quad\text{ for all }l\in\overline{\Lambda}.$ (3.5)
Indeed, let us assume (3.4). In this case, we take
$\bar{z}=-1+\min_{l\in\overline{\Lambda}}\left(x_{l}(T_{3})-\bar{\varphi}_{2}(l)\right).$
Let $T_{4}>T_{3}$ satisfy
$\begin{split}&\frac{3\alpha|\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3})|(T_{4}-T_{3})^{2}+6.75\sqrt{\alpha\beta}|\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3})|^{2}(T_{4}-T_{3})}{(T_{4}-T_{3})^{3}}\\ &+\frac{3.375\beta|\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3})|^{3}+6|\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3})|(T_{4}+T_{3})}{(T_{4}-T_{3})^{3}}<M+\Delta\bar{\varphi}_{2}(l)\end{split}$
for all $l\in\overline{\Lambda}$. Denoting by
$v_{l}(t)=\bar{v}+\frac{6(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))}{(T_{4}-T_{3})^{3}}(t-T_{3})(T_{4}-t)\leq\bar{v}\quad\text{ for }t\in[T_{3},T_{4}],$
we obtain
$\dot{v}_{l}(t)=\frac{6(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))}{(T_{4}-T_{3})^{3}}(T_{3}+T_{4}-2t)\geq 0\quad\text{ for }t\in[(T_{3}+T_{4})/2,T_{4}].$
Then, for $t\in[(T_{3}+T_{4})/2,T_{4}]$ which is close to $T_{4}$,
[…]
Therefore, we can conclude that
$M\leq u_{l}(t)<M\quad\text{ for all }t\in[T_{3},T_{4}].$
Thus, Theorem 1.2 still holds if we replace (1.5) by either (3.4) or (3.5).
3.2. Proof of Theorem 1.3
Sketch of the proof of Theorem 1.3.
Since this is a by-product of Theorems 1.11.2, we only give a sketch of the proof here.
By the proof of Theorem 1.1, we know that given any initial conditions, we can drive the system to a stationary flock at time $t=T_{1}$ satisfying
$\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(T_{1})-x_{l}(T_{1})}{h^{2}}\right|\leq\epsilon$ (3.6)
for all $l\in\overline{\Lambda}$.
We then use the first step in the proof of Theorem 1.2 to drive the system to a stationary flock where all particles are at the same position at time $T_{5}>T_{1}$. Afterwards, we use the second step in the proof of Theorem 1.2 to drive the system to the desired stationary flock at $T_{6}>T_{5}$. That is, for $l\in\overline{\Lambda}$ and $t\geq T_{6}$,
$\begin{cases}x_{l}(t)=\bar{\varphi}_{3}(l)+a,\\ v_{l}(t)=\bar{v}=0.\end{cases}$
4. Connection to optimal control and Hamilton–Jacobi equations
The goal of the control strategy above is to drive the system to a flock in a finite amount of time. Although the strategy strives to do this efficiently, it is by no means optimal. A related problem is to find the minimal time it takes to start from a given state and reach a flock with a control that has magnitude less than $M$. There is a rich theory behind this kind of optimal control problem. Here, we connect this optimal problem to a Hamilton-Jacobi problem.
The first step is to consider an ODE that does not have an interaction term like the one in (1.3). Thus, we write $X_{1}=(x_{l})_{l\in\overline{\Lambda}}$ and $X_{2}=(v_{l})_{l\in\overline{\Lambda}}$ as vectors in $\mathbb{R}^{D^{n}}$. Denote by $X=(X_{1},X_{2})\in\mathbb{R}^{2D^{n}}$. Then, (1.3) can be written as
$\dot{X}=\tilde{b}(X,W)=b(X)+(0,W),$
with the given initial condition
$X(0)=\left((w_{0}(l))_{l\in\overline{\Lambda}},(w_{1}(l))_{l\in\overline{\Lambda}}\right).$
Here, the vector field $b$ is
$b(X)=\left(X_{2},\left(\sum_{l^{\prime}\sim l}\frac{(X_{1})_{l^{\prime}}-(X_{1})_{l}}{h^{2}}+(\alpha-\beta|(X_{2})_{l}|^{2})(X_{2})_{l}\right)_{l\in\overline{\Lambda}}\right).$
In this framework, $W$ is the control, and the admissible set is
$\overline{\mathcal{A}}_{M}=\left\{W\in\mathbb{R}^{D^{n}}\,:\,|W_{l}|=|u_{l}|\leq M\qquad\text{ for }l\in\overline{\Lambda}\right\}.$
Basically, we can think of $X$ as the master state since it contains all information, and the ODE for $X$ is then local in nature.
Without loss of generality, we are only concerned with moving flocks here. For a given state $X$, we are interested in the optimal time it takes to reach a moving flock, that is,
$\Gamma=\left\{X\;:\,\left|\sum_{l^{\prime}\sim l}\frac{(X_{1})_{l^{\prime}}-(X_{1})_{l}}{h^{2}}\right|\leq M,(X_{2})_{l}=-\sqrt{\alpha/\beta}\quad\text{ for }l\in\overline{\Lambda}\right\}\\ \bigcup\left\{X\;:\,\left|\sum_{l^{\prime}\sim l}\frac{(X_{1})_{l^{\prime}}-(X_{1})_{l}}{h^{2}}\right|\leq M,(X_{2})_{l}=\sqrt{\alpha/\beta}\quad\text{ for }l\in\overline{\Lambda}\right\}.$
Let $U(X)$ be the minimal time it takes to start from state $X$ to reach $\Gamma$. Then, for $X=(X_{1},X_{2})\in\mathbb{R}^{2D^{n}}$ and $P=(P_{1},P_{2})\in\mathbb{R}^{2D^{n}}$, the corresponding Hamiltonian is
$H(X,P)=\sup_{W\in\overline{\mathcal{A}}_{M}}\left(-\tilde{b}(X,W)\cdot P-1\right)=-b(X)\cdot P+M|P_{2}|_{1}-1.$
Here, $|\cdot|_{1}$ is the usual $1$-norm in $\mathbb{R}^{D^{n}}$. Therefore, $U$ satisfies the following eikonal-type equation in the viscosity sense
$\begin{cases}-b(X)\cdot D_{X}U(X)+M|D_{X_{2}}U(X)|_{1}=1\qquad&\text{ for }X\in\mathbb{R}^{2D^{n}},\\ U(X)=0\qquad&\text{ for }X\in\Gamma.\end{cases}$ (4.1)
See [32, Chapter 2]. We note that this is a very complicated PDE posed in high dimensions, as the vector field $b$ contains much information about the dynamics. The value function $U(X)$ gives us the minimal time to reach a moving flock starting from $X\in\mathbb{R}^{2D^{n}}$, but does not give us directly an optimal strategy to reach a moving flock. By Theorem 1.1, we have that $U$ is finite. As the Hamiltonian $H$ here is noncoercive in $P$, the regularity of the viscosity solution $U$ is not yet known.
The advantage of Theorems 1.11.2 is that they provide more or less explicit strategies to reach a moving flock in finite time. Nevertheless, it is important to emphasize that an optimal control framework is hidden behind this problem.
Equation (4.1) is interesting in its own right and merits further study. Specifically, the regularity of the viscosity solution $U$, the properties of the optimal trajectories (strategies) and possible approximations are of particular interest.
5. Conclusion
This work is the first in controlling Klein-Gordon chains and lattices. Due to the complexity of the nonlinear ODE system under consideration, we focus here on a periodic domain and on controls applied at every node.
Our analysis provides insight into the system’s dynamics. Building on this understanding, we plan to investigate the problem with sparse controls, where the evolution of neighboring nodes will be used to influence nodes for which $u_{l}\equiv 0$.
In future work, we also aim to study the problem without assuming periodicity. This leads to an infinite-dimensional ODE system, for which the Lyapunov functional developed in this paper is not well defined. Our intended strategy is to first steer the infinite-dimensional system toward a periodic regime and then apply the results obtained here. Thus, the present paper constitutes a crucial first step in this broader research program.